### v2

# Dataset Distillation as Pushforward Optimal Quantization

Abstract
Dataset distillation aims to find a synthetic training set such that training on the synthetic data achieves similar performance to training on real data, with orders of magnitude less computational requirements. Existing methods can be broadly categorized as either bi-level optimization problems that have neural network training heuristics as the lower level problem, or disentangled methods that bypass the bi-level optimization by matching distributions of data. The latter method has the major advantages of speed and scalability in terms of size of both training and distilled datasets. We demonstrate that when equipped with an encoder-decoder structure, the empirically successful disentangled methods can be reformulated as an optimal quantization problem, where a finite set of points is found to approximate the underlying probability measure by minimizing the expected projection distance. In particular, we link existing disentangled dataset distillation methods to the classical optimal quantization and Wasserstein barycenter problems, demonstrating consistency of distilled datasets for diffusion-based generative priors. We propose Dataset Distillation by Optimal Quantization, based on clustering in a latent space. Compared to the previous SOTA method D4M, we achieve better performance and inter-model generalization on the ImageNet-1K dataset with trivial additional computation, and SOTA performance in higher image-per-class settings. Using the distilled noise initializations in a stronger diffusion transformer model, we obtain SOTA distillation performance on ImageNet-1K and its subsets, outperforming diffusion guidance methods.
1 Introduction
Training powerful neural networks requires a large amount of data, and thus induces high computational requirements. Dataset distillation (DD) targets this computational difficulty by changing the data, as opposed to other parts of training such as optimization or architecture [53]. The DD objective consists of finding a synthetic training set, such that training a neural network on the synthetic data yields similar performance.
There are several closely related notions of reducing computational load when training new models on datasets. Core-set methods find a subset of training data (as opposed to synthetic data) such that a model trained on said subset will have similar performance [30, 11]. Model distillation, sometimes known as knowledge distillation, aims to train a smaller (student) model that predicts the output of a pre-trained (teacher) model [12, 40]. Importance sampling methods accelerate training by weighting training data, finding examples that are more influential for training [36]. For more detailed surveys on dataset distillation methods and techniques, we refer to [57, 44].
1.1 Bi-level formulation of dataset distillation
We first present the DD problem statement as given in [44]. Denote a training set by $\mathcal{T}$ (more generally, distribution of training data), and the expected and empirical risks (test and training loss) by $\mathcal{R}$ and $\mathcal{L}$ respectively, evaluated for a network parameter $\Phi$. Given a learning algorithm $\Phi=\Phi_{S}$, which takes a dataset $S$ and returns an $\mathcal{L}$-optimal parameter, the goal of DD is to find a synthetic dataset $\mathcal{S}$ (of given size) minimizing the test loss discrepancy:
$\mathcal{S}=\operatorname*{arg\,min}_{S}\left|\mathcal{R}(\Phi_{S})-\mathcal{R}(\Phi_{\mathcal{T}})\right|.$ (1)
This formulation is computationally intractable. Approximations include replacing the minimum discrepancy objective with maximum test performance, replacing the learning algorithm $\Phi$ with an inner neural network optimization problem, and solving the outer minimization problem using gradient methods. Common heuristic relaxations to the bi-level formulation 1 include meta-learning [53, 10], distribution matching [58], and trajectory matching [5]. Other methods include neural feature matching [59, 28] and the corresponding neural tangent kernel methods [33, 32], representative matching [26], and group robustness [52]. For better scaling, [5, 31] consider using generative priors such as GANs to generate more visually coherent images, replacing the need for optimization with a neural network inversion task and increasing performance.
While the bi-level formulation follows naturally from the qualitative problem statement of dataset distillation, there are two main drawbacks, namely computational complexity and model architecture dependence. This limits the applicability to large-scale datasets and imposes additional requirements when using the distilled data.
The first prototypical example of using a generative model to alleviate the aforementioned drawbacks of the bi-level formulation is by “factorizing” the high-dimensional image-space optimization. This consists of jointly training generative “hallucinator” networks to generate samples from a common set of learnable “base” points [25]. This was able to increase cross-architecture performance when compared to previous image-space methods.
1.2 Disentangled methods
The dimensionality of the underlying bi-level optimization problems limits the applicability on large scale datasets, which are particularly useful for computationally limited applications. For example, the ImageNet-1K dataset consists of 1.2M training images, totalling over 120GB of memory [9]. The full dataset ImageNet-21K consists of over 14M images and takes up around 1.2TB of memory, which is generally infeasible to train expert models on, and makes backpropagation through network training steps impossible.
[56] is the first work to “disentangle” the bi-level optimization framework into three separate problems, named Squeeze, Recover and Relabel (SRe2L). In particular, the inner neural network optimization problem is replaced with matching statistics of batch-normalization layers. Curriculum Data Augmentation (CDA) uses adaptive training to get more performance [55]. [24] considers optimizing images such that neural network features are close to Wasserstein barycenters of the training image features. [51] proposes Realistic Diverse and Efficient Dataset Distillation (RDED), which replaces the latent clustering objective with a patch-based adversarial objective. [50] considers clustering directly in the latent space of a latent diffusion model (LDM) [43], named Dataset Distillation via Disentangled Diffusion Model (D4M). This avoids backpropagation when distilling and has constant memory usage with respect to images per class (IPC), a direct advantage over the linear memory scaling of optimization-based methods. Recent state-of-the-art methods consider fine-tuning diffusion models to have better statistics [14], and guiding the diffusion to maximize the influence of the distilled points [6], similarly to active learning.
While dataset distillation has had extensive experimental effort, few proper theoretical justifications or computable formalizations exist in the literature. [44] proposes a high-level formulation based on minimizing the difference in test loss between learning on the full dataset versus the synthetic dataset. [22] considers dataset distillation as dependent on the desired inference task (typically classification with cross-entropy loss for image data). This allows them to interpret trajectory-matching as a mean-field control problem, and find analogous procedures for non-imaging examples like finding collocation points for training physics informed neural networks. However, they do not consider the problem of consistency or convergence of the distilled datasets, which we aim to address in this work.
We interpret dataset distillation as the task of finding a surrogate dataset that minimizes the Wasserstein distance with the original training dataset. As a consequence for training neural networks, the gradient steps for a well-distilled dataset will be similar to the steps for the training dataset. This generalizes the objectives of the bi-level approaches to allow for different network architectures.
1.3 Contributions
We summarize the contributions of this work as follows.
  1. 1.
    Disentangled dataset distillation can be theoretically justified using classical notions of optimal quantization and Wasserstein distance. Motivated by the empirical usage of clustering in latent spaces, we show in Theorem˜1 that optimal quantizations induce convergent approximations of gradients of population risk, and furthermore the rate is given by $\mathcal{O}(K^{-1/d})$, where $d$ is the dimension of the latent space and $K$ is the number of quantization points. This motivates the usage of a low-dimensional latent space.
  2. 2.
    From the aforementioned classical literature, we propose Dataset Distillation by Optimal Quantization (DDOQ) in Algorithm˜2, based on clustering in a low-dimensional latent space. This is related to a recent disentangled method D4M. Our proposed method has a smaller Wasserstein distance between the distilled latent points and latent data distribution. This theoretically results in network gradients that more closely approximate those from training on the full training dataset.
  3. 3.
    We algorithmically compare our proposed method with D4M and various baselines on the ImageNet-1K dataset using the same generative diffusion model backbone, demonstrating significantly better classification accuracy at varied IPC budgets and better cross-architecture generalization. Moreover, using the stronger diffusion transformer backbone, we provide a central comparison with SOTA disentangled and diffusion-based distillation methods, yielding competitive or better results than existing state-of-the-art on ImageNet-1K and some subsets.
This paper is structured as follows. Section˜2 covers related background and convergence results, including optimal quantization and diffusion models. Theorem˜1 in Section˜3 demonstrates consistency of the optimal quantizers when passed through diffusion-based generative priors to the image space, and proposes a simple extension of the D4M method by adding automatically-learned weights in the prototyping phase. We further fully explain the data distillation pipeline in a sequential and systematic manner from distillation, to training new models using the computed weights. Section˜4 contains experiments of our proposed method against the previously SOTA D4M method as well as other recent SOTA baselines on the large-scale ImageNet-1K dataset.
2 Background
This section covers the necessary background and concepts required for the theoretical interpretation of DD. Section˜2.1 first covers necessary notation. Section˜2.2 introduces the optimal quantization problem, the concept of quadratic distortion, and links it to Wasserstein optimality. Section˜2.3 provides a short exposition of score-based diffusion models, covering well-posedness and classical divergence bounds.
2.1 Definitions and Notation
Define $\mathcal{P}_{2}(\mathbb{R}^{d})$ to be the set of probability measures on $\mathbb{R}^{d}$ with finite second moment, not necessarily admitting a density with respect to the Lebesgue measure. We use ${\mathcal{W}}_{2}$ to denote the Wasserstein-2 distance between two probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ [45], defined as
${\mathcal{W}}_{2}(\mu,\nu)=\left(\inf_{\gamma\in\Gamma(\mu,\nu)}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\ \mathrm{d}\gamma(x,y)\right)^{1/2},$ (2)
where $\Gamma(\mu,\nu)$ denotes the set of all couplings, i.e. joint probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$.
2.2 Optimal Quantization
For a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with finite second moments, an optimal quantization (or vector quantization) at level $K$ is a set of points $\{x_{1},...,x_{K}\}\subset\mathbb{R}^{d}$ such that the $\mu$-averaged Euclidean distance to the quantized points is minimal. This can be formulated as the minimizer of the (quadratic) distortion, defined as follows.
Definition 1 (Quadratic distortion).
For a quantization grid $(x_{1},...,x_{K})\in(\mathbb{R}^{d})^{K}$, the corresponding Voronoi cells are
$C_{i}=\{y\in\mathbb{R}^{d}\mid\|y-x_{i}\|=\min_{j}\|y-x_{j}\|\},\quad i=1,...,K.$ (3)
Given a measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the (quadratic) distortion function ${\mathcal{G}}={\mathcal{G}}_{\mu}$ takes a tuple of points $(x_{1},...,x_{K})$ and outputs the average squared distance to the set:
${\mathcal{G}}:(x_{1},...,x_{K})\mapsto\int_{\mathbb{R}^{d}}\min_{i}\|x-x_{i}\|^{2}\,\mu(\mathrm{d}x)=\mathbb{E}_{X\sim\mu}[\min_{i}\|X-x_{i}\|^{2}].$ (4)
We will write ${\mathcal{G}}_{K,\mu}$ to mean the distortion function at level $K$, i.e. with domain $(\mathbb{R}^{d})^{K}$, and drop the subscripts where it is clear. An optimal quantization is a minimizer of $\mathcal{G}$.
Note that the assumption that $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ has finite second moments implies that the quadratic distortion is finite for any set of points, and that an optimal quantization exists [34]. We note that the optimal quantization weights are uniquely determined by the quantization points. This gives equivalence of the distortion minimization problem to the Wasserstein minimization problem, when restricted to measures of finite support [34].
Proposition 1.
Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$. Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial C_{i})=0$. Then the measure $\nu$ that minimizes the Wasserstein-2 distance 2 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$. Moreover, the optimal coupling is given by the projection onto the centroids.
For a given quantization $\mathbf{x}$} or quantization level $K$, the quadratic distortion satisfies respectively:
${\mathcal{G}}({\mathbf{x}})^{1/2}=\inf\left\{{\mathcal{W}}_{2}(\nu,\mu)\mid\text{probability measures }\nu,\,\operatorname{supp}\nu\subset{\mathbf{x}}\right\};$ (5)
$\operatorname*{arg\,min}_{{\mathbf{x}},\,|{\mathbf{x}}|=K}{\mathcal{G}}({\mathbf{x}})=\operatorname*{arg\,min}\left\{{\mathcal{W}}_{2}(\nu,\mu)\mid\text{probability measures }\nu,\,|\operatorname{supp}\nu|\leq K\ \right\}.$ (6)
In other words, minimizing the quadratic distortion is equivalent to finding a Wasserstein-2 optimal approximation with a $K$-finitely supported (probability) measure.
Remark 1.
The case where the approximating measure is a uniform Dirac mixture is called the Wasserstein barycenter problem [8]. The Wasserstein error of the Wasserstein barycenter is higher than that of the optimal quantization, but it admits an easily computable dual representation.
Optimal quantizers converge in Wasserstein distance at a rate $\Theta(K^{-1/d})$ to their respective distributions as the number of particles $K$ increases [13]. Moreover, any limit point of the optimal quantizers is an optimal quantizer of the limiting distribution [41]. Moreover, it can be shown that if a sequence of distributions converges in Wasserstein distance $\mu_{n}\rightarrow\mu$, then so do the errors in quantization for a fixed quantization level [27, Thm. 4].
The quantizer can also be shown to have nice approximation properties when taking expectations of functions. A prototypical example for DD would have $f$ be the gradient of a neural network with respect to some loss function. This implies that a data distribution $\mu$ and its quantization $\nu$ induce similar training dynamics.
Proposition 2.
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be an $L$-Lipschitz function. For a probability measure $\mu\in{\mathcal{P}}_{2}(\mathbb{R}^{d})$ that assigns no mass to hyperplanes, and a quantization ${\mathbf{x}}=(x_{1},...,x_{k})$, let $\nu=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$ be the corresponding Wasserstein-optimal measure with support in $\mathbf{x}$}, as in Proposition˜1. The difference between the population risk $\mathbb{E}_{\mu}[f]$ and the weighted empirical risk $\mathbb{E}_{\nu}[f]$ is bounded as
$\mathbb{E}_{\mu}[f]-\mathbb{E}_{\nu}[f]\leq L{\mathcal{G}}({\mathbf{x}})^{1/2}.$ (7)
Proof.
Deferred to Section˜C.1. ∎
2.2.1 Solving the optimal quantization problem
To find an optimal quantizer, we now introduce a commonly used quantization algorithm: the competitive learning vector quantization (CLVQ) algorithm [1]. This is derived directly from gradient descent on the quadratic distortion, starting from an initialization in the convex hull of the support of $\mu$. For a given quantization, the gradient of the distortion has a representation in terms of $\mu$-centroids of the corresponding Voronoi cells, given explicitly in Section˜C.3.
The CLVQ algorithm is presented in Algorithm˜1. It consists of iterating: (i) sampling from $\mu$, (ii) computing the nearest cluster centroid, and (iii) updating the cluster centroid with weighted average. This algorithm is equivalent to the mini-batch $k$-means method when the stepsizes $\gamma_{i}$ are chosen to be the reciprocals of the number of points per-cluster [46, 37]. We can thus interpret mini-batch $k$-means as finding a local minima of the optimal quantization problem.
While this algorithm produces points ${\mathbf{x}}=(x_{1},...,x_{K})$, it remains to compute the associated weights approximating the measures of the Voronoi cells $\mu(C_{i})$ as in Proposition˜1. This can be done in an online manner within the same iterations [34].
Data: initial cluster centers $x_{1}^{(0)},...,x_{K}^{(0)}$, step-sizes $(\gamma_{i})_{i\geq 0}$, $i\leftarrow 0$
1 Initialize weights ${\mathbf{w}}=(w_{1},...,w_{K})=(1/K,...,1/K)$;
2 while not converged do
3   Sample $X_{i}\sim\mu$;
4   Select “winner” $k_{\textrm{win}}\in\operatorname*{arg\,min}_{1\leq k\leq K}\|X_{i}-x_{k}^{(i)}\|$;
5   Update $x_{k}^{(i+1)}\leftarrow(1-\gamma_{i})x_{k}^{(i)}+\gamma_{i}X_{i}$ if $k=k_{\mathrm{win}}$, otherwise $x_{k}^{(i+1)}\leftarrow x_{k}^{(i)}$ ;
6   Update weights $w_{k}\leftarrow(1-\gamma_{i})w_{k}+\gamma_{i}\mathbf{1}_{k=k_{\mathrm{win}}}$;
7 $i\leftarrow i+1$;
8
9 end while
Result: quantization $\nu_{K}=\sum_{k=1}^{K}w_{i}\delta(x_{k}^{*})$
Algorithm 1 CLVQ
Proposition 3 ([3, Prop. 7]).
Assume that the measure $\mu\in{\mathcal{P}}_{2+\eta}(\mathbb{R}^{d})$ for some $\eta>0$, and that it assigns no mass to hyperplanes. Assume further that the grids ${\mathbf{x}}^{(t)}$ produced by CLVQ Algorithm˜1 converge to a stationary grid $\mathbf{x}$$^{*}$, i.e. $\nabla{\mathcal{G}}({\mathbf{x}}^{*})=0$, and that the step-sizes satisfy $\sum\gamma_{k}=+\infty$ and $\sum\gamma_{k}^{1+\delta}<\infty$ for some $\delta>0$. Then,
  1. 1.
    The companion weights $w_{k}$ converge almost surely to the limiting weights $\mu(C_{k}^{*})$;
  2. 2.
    The moving average of the empirical quadratic distortion converges to the limiting distortion:
    $\frac{1}{t}\sum_{k=1}^{t}\min_{1\leq i\leq K}\|X_{k}-x_{i}^{(k)}\|^{2}\rightarrow{\mathcal{G}}({\mathbf{x}}^{*}).$
Using these weights, we target the problem of approximating optimal quantizers, rather than the Wasserstein barycenter problem. The addition of the weights reduces the distortion in the latent space. In the following section, we show that this reduced distortion carries over to the image space, which leads to better training fidelity as given using Proposition˜2.
Remark 2.
Convergence to an optimal quantizer is only guaranteed in the case of a log-concave distribution in one dimension [21, 27]. However, in higher dimensions, convergence to a (not necessarily optimal) stationary grid is possible [35, 34].
2.3 Score-based diffusion
To connect the quantization error on the latent space with the quantization error in the image space, we need to consider properties of the latent-to-image process. In particular, we focus on score-based diffusion models, interpreted as discretizations of particular noising SDEs [49]. Consider the following SDE, where $W$ is a standard Wiener process:
[…]
Then, where the density of $x$ at time $t>0$ is given by $p_{t}$, the reverse of the diffusion process is given by the reverse-time SDE [16],
[…]
where $\bar{W}$ is the standard Wiener process running in reverse time from time $T$ to 0. For an increasing noising schedule $\sigma(t)\in[0,+\infty)$ or noise-scale $\beta(t)\in[0,1)$, the variance-exploding SDE (VESDE, or Brownian motion) is given by
[…]
and the variance-preserving SDE (VPSDE, or Ornstein–Uhlenbeck process) is given by
[…]
The VPSDE and VESDE are commonly used when training diffusion models due to their simple discretization. Indeed, they both arise as continuous limits of earlier proposed diffusion models. Specifically, VPSDE corresponds to denoising score matching with Langevin dynamics [48], and VESDE to denoising diffusion probabilistic models [47, 17].
Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with bounded support. [39] demonstrates that the backward SDE processes are indeed diffusions for the VPSDE and VESDE. Moreover, the diffusion satisfies a data-processing inequality, showing that the $f$-divergence after backwards diffusion is at most the $f$-divergence at marginal time $T$. We review this in Appendix˜D.
Convergence in $f$-divergence necessarily requires absolute continuity of the compared marginal with respect to the underlying diffused distribution, which is equivalent to continuity with respect to the Lebesgue measure by the Hörmander condition. This rules out singular initializations such as empirical measures, as one might want to use for dataset distillation. To combat this, we use weak convergence with appropriately chosen test functions.
The goal in question: given Wasserstein-2 convergence of the marginals $\nu_{T}^{(k)}\rightarrow\mu_{T}$, we wish to derive a bound on expectations $\mathbb{E}_{\nu_{\delta}^{(k)}}[f]\xrightarrow{?}\mathbb{E}_{\mu_{\delta}}[f]$, for some fixed $\delta\in(0,T)$ and $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying some regularity conditions. Such a bound would directly link to training neural networks with surrogate data, e.g. by taking $f$ to be the gradient of a loss function with respect to some network parameters.
Remark 3.
Having $\delta=0$ may not be well defined because of non-smoothness and blowup of the score at time 0 for singular measures [39, 54].
3 Dataset Distillation as Optimal Quantization
Our main result Theorem˜1 gives consistency of dataset distillation for a score-based diffusion prior in the image space. It transforms convergence of the terminal distribution into convergence of the backwards-diffused distributions at small time. Later, we use this in Section˜3.1 to present DDOQ as a modification of the D4M method. By simply changing the clustering objective from a Wasserstein barycenter to an optimal quantization by adding weights, we can effectively reduce the Wasserstein distance to the data distribution.
Theorem 1.
Consider the VESDE/Brownian motion
[…]
or the VPSDE/Ornstein–Uhlenbeck process
[…]
For any initial data distribution $\mu\in{\mathcal{P}}_{2}(\mathbb{R}^{d})$ with compact support bounded by $R>0$, the backwards diffusion process is well posed. Suppose that there are two distributions $\mu_{T},\nu_{T}$ at time $T$ that undergo the reverse diffusion process (with fixed initial reference measure $\mu$) up to time $t=\delta\in(0,T)$ to produce distributions $\mu_{\delta},\nu_{\delta}$. There exists a (universal explicit) constant $C=C(\delta,T,R,d)\in(0,+\infty)$ such that if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ is an $L$-Lipschitz function, then the difference in expectation satisfies
$\|\mathbb{E}_{\mu_{\delta}}[f]-\mathbb{E}_{\nu_{\delta}}[f]\|\leq CL{\mathcal{W}}_{2}(\mu_{T},\nu_{T}).$ (14)
In dataset distillation terms, $f$ will typically be replaced by the gradient of a loss function. The above result suggests that a distilled image dataset can be given by passing a distilled latent dataset through the generative reverse SDE process. Moreover, when training a neural network on a distilled dataset given by optimal quantization, the gradients on the distilled dataset and full training dataset at each step will automatically be similar. This bypasses the heuristics needed in bi-level DD formulations, and avoids fine-tuning or generation-time guidance of the diffusion models.
Theorem˜1 combined with the asymptotic rates of [13] gives convergence rates as the number of quantization points increases. We note that this can be further be combined with the convergence of optimal quantization rates of empirical measures such as Theorem˜3 in Section˜C.2.
Corollary 1.
Suppose $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ has compact support and is diffused through either the Brownian motion or Ornstein–Uhlenbeck process up to time $T$ to produce marginal $\mu_{T}$. Let $\nu^{(K)}_{T}$ be optimal quantizers of $\mu_{T}$ at level $K$ for $K\in\mathbb{N}$. For fixed $\delta\in(0,T)$, let $\nu^{(K)}_{\delta}$ denote the corresponding backwards diffusion at time $T-\delta$. Then, for any $L$-Lipschitz function $f$ and as $K\rightarrow\infty$,
$\|\mathbb{E}_{\mu_{\delta}}[f]-\mathbb{E}_{\nu^{(K)}_{\delta}}[f]\|=L\mathcal{O}(K^{-1/d}).$ (15)
In particular, as the number of points increases, we have convergence to the underlying data distribution in image space, giving consistency.
Data: Training data and labels $({\mathcal{T}},{\mathcal{L}})$, pre-trained encoder-decoder pair $({\mathcal{E}},{\mathcal{D}})$, text encoder $\tau$, latent diffusion model ${\mathcal{U}}_{t}$, target IPC count $K$
/* Step 1: encode */
1 Initialize latent points $Z={\mathcal{E}}({\mathcal{T}})$;
/* Step 2: cluster */
2 Compute and save $k$-means cluster centers $z_{k}^{(L)}$ and cluster counts $v_{k}^{(L)}$, $k=1,...,K,\,L\in{\mathcal{L}}$;
3 Compute weights $w_{k}^{(L)}\leftarrow v_{k}^{(L)}/\sum_{j}v_{j}^{(L)}$;
/* Step 3: decode */
4 Compute class embeddings $\mathrm{emb}=\tau(L),\,L\in{\mathcal{L}}$;
5 Compute and save class distilled images ${\mathcal{S}}_{L}=\{x^{(L)}_{k}={\mathcal{D}}\circ{\mathcal{U}}_{t}(z^{(L)}_{k},\mathrm{emb})\mid k=1,...,K\}$;
Result: Distilled images ${\mathcal{S}}=\bigcup_{L\in{\mathcal{L}}}{\mathcal{S}}_{L}$
6 To train a new network $f_{\theta}$:
Data: Labels $y$ for training data $x\in\mathcal{S}$, loss function $\ell:(x,y,\theta)\to\mathbb{R}$
/* Step 4: Train new model (Validation) */
Train network using loss function $\min_{\theta}\sum_{(x,y,w)}w\cdot\ell(x,y,\theta)$
Algorithm 2 Dataset Distillation by Optimal Quantization (DDOQ)
3.1 Proposed method: Dataset Distillation by Optimal Quantization
We have seen that clustering gives a consistent approximation to the latent distribution. Moreover, using the generative diffusion model, this then gives a consistent approximation in the image space. Now we present the proposed algorithm, Dataset Distillation by Optimal Quantization (DDOQ). We detail the steps in text below and summarize in Algorithm˜2.
Suppose we are given an LDM, i.e. a (conditional) encoder-decoder pair and a diffusion-based model on the latent space [43]. Let $K$ be the target number of images per class. Constructing the distilled dataset consists of the following steps in sequence.
  1. 1.
    Encoding. We use the encoder of the LDM to map training samples from the image space to the latent space, giving empirical samples from the SDE marginal $\mu_{T}$. This is done per-class in the context of classification.
  2. 2.
    Clustering. We then use the CLVQ (mini-batch $k$-means) algorithm on these samples to compute the $K$ centroids and corresponding weights, as in Proposition˜3. This gives an empirical distribution $\nu_{T}^{(K)}$ that approximates $\mu_{T}$. Equivalently, it consists of tuples of points and weights $(x_{i},w_{i})_{i=1}^{K}$, such that $\nu_{T}^{(K)}=\sum_{i}w_{i}\delta(x_{i})\approx\mu_{T}$.
  3. 3.
    Decoding/Image synthesis. Given the clustered centroids in the latent space, the generative part of the LDM is employed to reconstruct images. This comprises the image component of the distilled dataset. The weights of the latent points are assigned directly to the weights of the corresponding generated image.
  4. 4.
    Training new models. Training a network $f_{\theta}$ from scratch requires: distilled images $x$, some corresponding labels $y$, and the weights $w$ to each image. For a loss function $\ell(x,y,\theta)$ such as cross-entropy or KL divergence, the loss for each sample is weighted by $w$. The complete loss function is given by
    $\min_{\theta}\sum_{(x,y,w)}w\cdot\ell(x,y,\theta).$
We note that there is flexibility in the choice of label $y$ when training a new model. For example, the soft-label synthesis of D4M or RDED employs another pre-trained classifier $\Psi$, such as a ResNet. Using this, the soft-labels are given by $y=\Psi(x)$, as opposed to one-hot encodings of the corresponding classes.
4 Experiments
We compare with two different latent diffusion architectures, namely the original latent diffusion model utilizing UNets [43], then the stronger diffusion transformer (DiT) architecture [38]. We first use the LDM to compare the pure algorithmic differences of DDOQ with D4M and related baselines. Then, we use DiT to demonstrate the potential of DDOQ compared to newer state-of-the-art baselines.
4.1 UNet backbone
Figure 1: Example distilled images of the “jeep” class in ImageNet-1K along with their $k$-means weights below. There are little to no qualitative features that can be used to differentiate the low and high weighted images, mainly due to the high fidelity of the diffusion model. However, the weights are indicative of the distribution of the training data in the latent space of the diffusion model.
Table 1: Comparison of top-1 classification performance on the ImageNet-1K dataset for baselines versus the proposed DDOQ method at various IPCs. We observe that DDOQ outperforms the previous SOTA method D4M, due to the addition of weights to the synthetic data. The maximum performance for all methods should be 69.8 as the soft labels are computed using a pre-trained ResNet-18 model. Figures for TESLA, SRe2L and D4M are taken directly from [50], RDED and CDA from [51, 55].
[…]
To validate the proposed DDOQ algorithm, we directly compare with the previous state-of-the-art methods D4M [50] and RDED [51] on the ImageNet-1K dataset. The latent diffusion model chosen for latent generation and image synthesis is the publicly available pre-trained Stable Diffusion V1.5 model. We report results for SRe2L [56], RDED [51], CDA [55] as given. For low IPC, we also report the TESLA method, which is a SOTA bi-level method based on MTT [7]. RDED achieves SOTA results for low IPC using its aggregated images and special training schedule, while CDA improves upon SRe2L using time-varying augmentation.
For consistency and a more direct comparison with previous methods, we use the pre-trained PyTorch ResNet-18 model to compute the soft labels, using the same protocol as [50]. After computing the soft labels using the pre-trained ResNet-18 model, we train new ResNet-18, ResNet-50 and ResNet-101 models to match the soft labels. The data augmentation is also identical, with the only difference being the addition of the weights to the training objective and minimal optimizer hyperparameter tuning for the new objective. We provide a direct comparison of the distilled data performance in Table˜1 for the IPCs $K\in\{10,50,100,200\}$. Our figures for DDOQ are averaged over five models trained on the same distilled data.
We observe that while RDED is very powerful in the low IPC setting due to the patch-based distilled images, which effectively gives the information of 4 (down-sampled) images in one training sample. However, the gap quickly reduces for IPC 50, getting outperformed by the clustering-based D4M and proposed DDOQ methods with more powerful models like ResNet-101. Results for IPC 100/200 are not available online for RDED and we omit them due to computational restriction.
The proposed DDOQ algorithm is uniformly better than D4M, with the most significant increase in the low IPC setting. Moreover, DDOQ surpasses all the compared SOTA disentangled DD methods for IPC 100 and 200. For a low number of quantization points, the gap in Wasserstein distance of the Wasserstein barycenter and the optimal quantizer to the data distribution may be large. As indicated in Theorem˜1, a lower Wasserstein distance means more faithful gradient computations on the synthetic data, which may explain the higher performance with minimal algorithmic change, as well as implicitly allowing for gradient matching as in existing bi-level methods.
Table 2: Generalization performance of D4M and the proposed DDOQ method for different soft-label teachers and student architectures at IPC 50. We observe that DDOQ has uniformly better cross-architecture generalization for convolutional teacher and student architectures, and slightly worse performance for student models using the transformer architecture Swin-T. Figures for D4M as reported in [50]. Bold indicates best method.
Teacher Network Student Network
ResNet-18 MobileNet-V2 EfficientNet-B0 Swin-T
ResNet-18 D4M 55.2 47.9 55.4 58.1
DDOQ 56.2 52.1 58.0 57.4
MobileNet-V2 D4M 47.6 42.9 49.8 58.9
DDOQ 47.7 45.6 52.5 56.3
Swin-T D4M 27.5 21.9 26.4 38.1
DDOQ 28.5 24.1 29.3 36.0
Table 3: Wasserstein-2 distances of the distilled latents compared to the encoded training data. We observe the weighting drops the Wasserstein-2 distance by around 20%.
Test class 0 1 2
IPC 10 D4M 49.67 51.14 47.16
DDOQ 42.97 42.71 39.14
IPC 50 D4M 47.34 49.10 45.28
DDOQ 40.41 41.01 37.53
Table˜2 shows the generalization performance of the distilled latent points. Different PyTorch pre-trained teacher networks are used to create the soft labels as in Step 3 of DDOQ, in particular MobileNet-V2 and Swin-T. We then evaluate the distilled images and soft labels with three convolutional architectures and the Swin-T transformer architecture. We observe that DDOQ is not only able to generalize to different model architectures, but also uniformly outperforms D4M on all convolutional student architectures. The slightly worse performance of DDOQ when using a transformer architecture for the student may be due to more precise hyperparameter tuning requirements.
To illustrate the weights, Figure˜1 plots ten example images from the “jeep” class when distilled using $K=10$ IPC. We observe that there is a very large variance in the weights $v_{k}^{(L)}\big{/}\sum_{j=1}^{K}v_{j}^{(L)}$}, indicating the presence of strong clustering in the latent space. Nonetheless, there is no qualitative evidence that the weights indicate “better or worse” training examples. Indeed, among other classes, there are synthetic images that do not lie in the appropriate class, but are still weighted highly.
To consider the effect of weighting the centroids, we consider the class-wise Wasserstein distance between the centroids and encoded training data, with and without the weights. Table˜3 plots the Wasserstein distance between the distilled latent points and the encoded training dataset, for the first three classes of ImageNet-1K. The latent points are clustered with uniform weights (D4M), and with variable weights (DDOQ). We observe that adding weights reduces the Wasserstein distance uniformly by around 20%. This leads to better gradient faithfulness when training a neural network.
4.2 DiT backbone
To compare the potential of DDOQ, we use a stronger generative model, namely the diffusion transformer [38]. This architecture achieves uniformly better sample quality compared to the LDM used in the previous section, in terms of Inception Score and Fréchet inception distance.
We compare with the SOTA DD methods based on diffusion guidance, namely Minimax [14] and Influence Guided Diffusion (IGD) [6], which both use DiT. These methods guide the decoding process using some fine-tuning or batch statistics. In contrast, DDOQ directly modifies the initialization in the latent space. In addition, we compare with samples directly generated with DiT from random initializations. We compare on ImageNet-1K, as well as the 10-class subsets ImageNette and ImageWoof. The baseline figures are taken from [6] which report higher numbers than [14].
Table˜4 demonstrates that we significantly outperform the diffusion-guidance based methods on the full ImageNet-1K, as well as DDOQ with the UNet backbone. Moreover, we are competitive-with or better than the baselines on the 10-class subsets, namely outperforming in the low IPC setting and in the more difficult ImageWoof subset.
Table 4: Comparisons on ImageWoof and ImageNette using the ResNetAP-10 architecture, and ImageNet-1K using ResNet-18. We observe that DDOQ-DiT is able to outperform the guided diffusion methods at low IPC and are competitive at higher IPC, namely outperforming the compared methods on ImageNet-1K. The maximum accuracies are 87.2, 94.6 and 69.8 for ImageWoof, ImageNette and ImageNet-1K respectively.
Dataset IPC DiT DiT-IGD Minimax Minimax-IGD DDOQ-DiT Full
[…]
5 Conclusion
This work proposes DDOQ, a dataset distillation method based on optimal quantization. Inspired by optimal quantization and Wasserstein theory, we theoretically demonstrate consistency of the distilled datasets in Theorem˜1 when using standard diffusion-based generative models to generate the synthetic data. We empirically demonstrate performance higher than previous state-of-the-art methods on a large-scale ImageNet-1K dataset, as well as in the low IPC setting of ImageNet subsets.
We have presented theoretical justification for one framework of dataset distillation. Future work could include sharper bounds in Theorem˜1 that exploit the sub-Gaussianity of the diffused distributions or manifold hypothesis, or investigating alternative choices of weight variance reduction. Other interesting directions could be relating the weightings of the synthetic data to the hardness of learning the data, such as in [20], further increasing the performance using different training regimes such as curriculum learning, or extending the theoretical analysis to inexact score matching. Possible empirical work could investigate correlations between the weights and influence [42].
Acknowledgments and Disclosure of Funding
This work was done while HYT was on placement in GSK.ai. HYT acknowledges support from GSK and the Masason Foundation.
Appendix A Details on heuristic bi-level distillation
  1. 1.
    (Meta-learning.) This uses the assumption that more data produces better results, replacing the risk matching objective with a risk minimization objective and removing the dependence on $\mathcal{T}$ [53, 10]. The learning algorithm $\Phi_{S}=\Phi_{S}(\theta_{0})$ with initial parameter $\theta_{0}$ is given by unrolling gradient steps $\theta_{t+1}=\theta_{t}-\eta\nabla\mathcal{L}_{S}(\theta_{t})$. The distilled dataset minimizes the training loss as trained on the distilled set up to $T$ iterations:
    $\operatorname*{arg\,min}_{S}\mathbb{E}_{\theta_{0}\sim p_{\theta}}\left[\mathcal{L}_{\mathcal{T}}(\theta_{T})\right].$ (16)
  2. 2.
    (Distribution matching.) Inspired by reproducing kernel Hilbert spaces and the empirical approximations using random neural network features, [58] proposes an alternate distillation objective that is independent of target loss. The minimization objective is
    $\operatorname*{arg\,min}_{\mathcal{S}}\mathbb{E}_{\theta\sim p_{\theta}}\|\frac{1}{|\mathcal{T}|}\sum_{x\in\mathcal{T}}\psi_{\theta}(x)-\frac{1}{|\mathcal{S}|}\sum_{x\in\mathcal{S}}\psi_{\theta}(x)\|^{2},$ (17)
    where $\psi_{\theta}$ are randomly initialized neural networks. This can be intuitively interpreted as matching the (first moment of) neural network features over the synthetic data.
  3. 3.
    (Trajectory matching.) For a fixed network architecture, this method aims to match the gradient information of the synthetic and true training datasets from different initializations [5]. The heuristic is that similar network parameters will give similar performance. In addition, the gradients are allowed to be accelerated by matching a small number of synthetic training steps with a large number of real data training steps: for $N\ll M$ steps, the Matching Training Trajectory (MTT) objective is (with abuse of notation):
    $\operatorname*{arg\,min}_{\mathcal{S}}\mathbb{E}_{\theta_{0}\sim p_{\theta}}\sum_{t=1}^{T-M}\frac{\|\theta_{t+N}^{\mathcal{S}}-\theta_{t+M}^{\mathcal{T}}\|^{2}}{\|\theta_{t+M}^{\mathcal{T}}-\theta_{t}^{\mathcal{T}}\|^{2}},$ (18)
    $\text{where }\ \theta_{t+1}^{\mathcal{T}}=\theta_{t}^{\mathcal{T}}-\eta\nabla\mathcal{L}_{\mathcal{T}}(\theta_{t}^{\mathcal{T}}),\,\theta_{t+i+1}^{\mathcal{S}}=\theta_{t+i}^{\mathcal{S}}-\eta\nabla\mathcal{L}_{\mathcal{S}}(\theta_{t+i}^{\mathcal{S}}),\,\theta_{t+0}^{\mathcal{S}}=\theta_{t}^{\mathcal{T}}.$
    Practical computational approximations include pre-training a large number of teacher networks from different initializations.
Appendix B Proof of Proposition 1
Proposition 4.
Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$. Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial(C_{i}))=0$. Then the measure $\nu$ that minimizes the Wasserstein-2 distance 2 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$. Moreover, the optimal coupling is given by the projection onto the centroids.
Proof.
For any coupling $\gamma\in\Gamma(\nu,\mu)$ between $\nu$ and $\mu$, we certainly have that
$\iint\|x-y\|^{2}\,\mathrm{d}{\gamma}(x,y)\geq\iint\operatorname{dist}({\mathbf{x}},y)^{2}\,\mathrm{d}{\gamma}(x,y)=\int\operatorname{dist}({\mathbf{x}},y)^{2}\mathrm{d}\mu(y).$
where the first inequality comes from definition of distance to a set ($\gamma$-a.s.) and the support condition on $\nu$, and the equality from the marginal property of couplings. The final term is attained when $\nu$ has the prescribed form: the coupling is $c(x_{i},y)=\mathbf{1}_{y\in C_{i}}$ and the corresponding transport map is projection onto the set $\mathbf{x}$} (defined $\mu$-a.s. from the null condition). This shows a lower bound of 2 that is attained. ∎
Appendix C Optimal quantization
C.1 Proof of Thm 2
Theorem 2.
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be an $L$-Lipschitz function. For a probability measure $\mu\in{\mathcal{P}}_{2}(\mathbb{R}^{d})$ that assigns no mass to hyperplanes, and a quantization ${\mathbf{x}}=(x_{1},...,x_{k})$, let $\nu=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$ be the corresponding Wasserstein-optimal measure with support in $\mathbf{x}$}, as in Proposition˜1. The difference between the population risk $\mathbb{E}_{\mu}[f]$ and the weighted empirical risk $\mathbb{E}_{\nu}[f]$ is bounded as
$|\mathbb{E}_{\mu}[f]-\mathbb{E}_{\nu}[f]|\leq L{\mathcal{G}}({\mathbf{x}})^{1/2}.$ (19)
Proof.
Since $\mu$ assigns no mass to hyperplanes, we may decompose into Voronoi cells.
[…]
$=L{\mathcal{G}}({\mathbf{x}})^{1/2}.$
The first equality comes from definition of $\nu$, and the inequalities from the Lipschitz condition and Hölder’s inequality respectively. ∎
C.2 Convergence of optimal quantization
The following result gives a convergence rate, assuming slightly higher regularity of the underlying probability measure.
Proposition 5 ([27]).
Let $\eta>0$, and suppose $\mu\in{\mathcal{P}}_{2+\eta}(\mathbb{R}^{d})$. There exists a universal constant $C_{d,\eta}\in(0,+\infty)$ such that for every quantization level,
$e^{*}_{K,\mu}\leq C_{d,\eta}\cdot\sigma_{2+\eta}(\mu)K^{-1/d},$ (20)
where $\sigma_{r}(\mu)=\min_{a\in\mathbb{R}^{d}}\mathbb{E}_{\mu}[\|x-a\|^{r}]^{1/r}$.
The following non-asymptotic result considers the convergence of optimal quantizers for a sequence of probability measures, converging in the Wasserstein sense.
Theorem 3 ([27, Thm. 4]).
Fix a quantization level $K\geq 1$. Let $\mu_{n},\mu\in{\mathcal{P}}_{2}(\mathbb{R}^{d})$ with support having at least $K$ points, such that ${\mathcal{W}}_{2}(\mu_{n},\mu)\rightarrow 0$ as $n\rightarrow\infty$. For each $n\in\mathbb{N}$, let ${\mathbf{x}}^{(n)}$ be an optimal quantizer of $\mu_{n}$. Then
${\mathcal{G}}_{K,\mu}({\mathbf{x}}^{(n)})-\inf_{{\mathbf{x}}}{\mathcal{G}}_{K,\mu}({\mathbf{x}})\leq 4e^{*}_{K,\mu}{\mathcal{W}}_{2}(\mu_{n},\mu)+4{\mathcal{W}}_{2}^{2}(\mu_{n},\mu),$ (21)
where $e^{*}_{K,\mu}=[\inf_{{\mathbf{x}}}{\mathcal{G}}_{K,\mu}({\mathbf{x}})]^{1/2}$ is the optimal error.
C.3 Gradient interpretation of CLVQ
Proposition 6 (Differentiability of distortion [34, Prop. 3.1]).
Let ${\mathbf{x}}=(x_{1},...,x_{K})\in(\mathbb{R}^{d})^{K}$ be such that the $x_{i}$ are pairwise distinct, and assume that $\mu(\partial C_{i})=0$. Then, the quadratic distortion is differentiable with derivative
$\nabla{\mathcal{G}}({\mathbf{x}})=\left(2\int_{C_{i}(x)}(x_{i}-\xi)\,\mu(\mathrm{d}\xi)\right)_{i=1,...,K},$ (22)
i.e., the gradient for quantization point $x_{i}$ points away from the $\mu$-centroid of its Voronoi cell.
The gradient step for some step-sizes $\gamma_{k}\in(0,1)$ reads
${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\gamma_{k}\nabla{\mathcal{G}}({\mathbf{x}}^{(k)}),\quad{\mathbf{x}}^{(0)}\in\mathrm{Hull}(\operatorname{supp}\mu)^{K},$ (23)
where $\mathrm{Hull}$ denotes the convex hull. Recalling that the gradient 22 is a $\mu$-expectation over $C_{i}(x)$, the corresponding stochastic Monte Carlo version of the above gradient descent yields the CLVQ algorithm 1:
${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\gamma_{k}\left(\mathbf{1}_{X_{k}\in C_{i}^{(k)}}x_{i}^{(k)}-X_{k}\right)_{1\leq i\leq K},\quad X_{k}\sim\mu.$ (24)
Note that the computation of this requires the ability to sample from $\mu$, as well as being able to compute the nearest neighbor of $X_{k}$ to the quantization set $\mathbf{x}$} (equivalent to the inclusion $X_{k}\in C_{i}^{(K)}$).
C.4 Lloyd I algorithm
Lloyd I. This consists of iteratively updating the centroids with the $\mu$-centroids, given by the $\mu$-average of the Voronoi cells. Clearly, if this algorithm converges, then the centroids are equal to the $\mu$-centroids and the grid is stationary. This is more commonly known as the $k$-means clustering algorithm, employed in common numerical software packages such as scikit-learn [37]. Convergence of the Lloyd-I algorithm can be found in e.g. [35].
Data: Probability distribution $\mu$ with finite first moment, initial cluster centers $x_{1}^{(0)},...,x_{K}^{(0)}$
1 $k\leftarrow 0$;
2 while not converged do
3   Compute Voronoi cells $C_{i}^{(k)}=\{y\in\mathbb{R}^{d}\mid\|x_{i}^{(k)}-y\|=\min_{j}\|x_{j}^{(k)}-y\|\}$;
4   Replace cluster centers with $\mu$-centroids $x_{i}^{(k+1)}\leftarrow(\int_{C_{i}^{(k)}}x\mu(\mathrm{d}x))/{\mu(C_{i}^{(k)})}$;
5 $k\leftarrow k+1$;
6
7 end while
Algorithm 3 Lloyd I ($k$-means)
Appendix D Background on well-posed diffusions
Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}(\mathbb{R}^{d})$, assumed to have bounded support. Then, by the Hörmander condition, the law of a random variable $(X_{t})_{t\geq 0}$ evolving under either the VPSDE or the VESDE will admit a density $p_{t}(x)$ for all $t>0$ with respect to the Lebesgue measure, that is smooth with respect to both $x$ and $t$ [18]. Using the following proposition, we have well-definedness of the backward SDE [2, 16].
Proposition 7.
For the forward SDE 8, assume that there exists some $K>0$ such that
  1. 1.
    $f(x,t),g(t)$ are measurable, and $f$ is uniformly Lipschitz: $\|f(x,t)-f(y,t)\|\leq K\|x-y\|$ for all $x,y\in\mathbb{R}^{d}$;
  2. 2.
    $\|f(x,t)\|+|g(t)|\leq K(1+\|x\|)$;
  3. 3.
    The solution $X_{t}$ of 8 has a $\mathcal{C}^{1}$ density $p_{t}(x)$ for all $t>0$, and
    $\int_{t_{0}}^{T}\int_{\|x\|<R}\|p_{t}\|^{2}+\|\nabla_{x}p_{t}(x)\|^{2}<+\infty\quad\forall t_{0}\in(0,T],\,R>0;$
  4. 4.
    The score $\nabla\log p_{t}(x)$ is locally Lipschitz on $(0,T]\times\mathbb{R}^{d}$.
Then the reverse process $X_{T-t}$ is a solution of the 9, and moreover, the solutions of 9 are unique in law.
Now given that the backwards SDE is indeed a diffusion, the data processing inequality uses the Markov property and states that the divergence after diffusion is less than the divergence before diffusion [23]. This is summarized in [39, Thm. 1].
Theorem 4.
Denote the initial data distribution by $\mu_{0}=\mu$, and let $\mu_{T}$ be the distribution of a random variable $X_{t}$ satisfying the forward SDE 8 on $[0,T]$. Assume the above assumptions, and let $Y_{t}$ satisfy the backwards SDE 9 on $[0,T]$ with terminal condition $Y_{T}\sim\nu_{T}$. Denote by $\mu_{t}$ and $\nu_{t}$ the marginal distributions at time $t\in[0,T]$ of $X_{t}$ and $Y_{t}$, and assume that $\nu_{T}\ll\mu_{T}$. Then:
  1. 1.
    The limit $Y_{0}\coloneqq\lim_{t\rightarrow 0^{+}}Y_{t}$ exists a.s., with distribution $\nu_{0}\ll\mu_{0}$.
  2. 2.
    For any $f$-divergence $D_{f}$,
    $D_{f}(\mu_{0},\nu_{0})\leq D_{f}(\mu_{T},\nu_{T})\quad\text{and}\quad D_{f}(\nu_{0},\mu_{0})\leq D_{f}(\nu_{T},\mu_{T}).$ (25)
This theorem shows that for an $f$-divergence, such as total variation distance or Kullback–Leibler divergence, convergence of the marginals at time $T$ implies convergence of the backwards-diffused marginals at time 0 (also at any $t<T$). However, this requires absolute continuity of the initial marginal distribution $\nu_{T}$ with respect to $\mu_{T}$, equivalently, w.r.t. Lebesgue measure. This precludes useful bounds for singular initializations of $\nu_{T}$, such as empirical measures.
Appendix E D4M algorithm
Data: Training data and labels $({\mathcal{T}},{\mathcal{L}})$, pre-trained encoder-decoder pair $({\mathcal{E}},{\mathcal{D}})$, text encoder $\tau$, latent diffusion model ${\mathcal{U}}_{t}$, target IPC count $K$
1 Initialize latent variables $Z={\mathcal{E}}({\mathcal{T}})$;
2 for label $L\in{\mathcal{L}}$ do
3   Initialize latent centroids $z^{k}_{L},\,k=1,...,K$;
4   Initialize update counts $v^{k}_{L}=1,\,k=1,...,K$;
 /* Compute prototypes with $k$-means */
5 for minibatch ${\mathbf{z}}\in Z|_{L}$ do
6    for $z\in{\mathbf{z}}$ do
7       $\hat{k}\leftarrow\operatorname*{arg\,min}_{k}\|z^{k}_{L}-z\|$;
       /* Update learning rate */
8       $v^{\hat{k}}_{L}\leftarrow v^{\hat{k}}_{L}+1$;
       /* Update centroid */
9       $z^{\hat{k}}_{L}\leftarrow(1-1/v^{\hat{k}}_{L})z^{\hat{k}}_{L}+(1/v^{\hat{k}}_{L})z$;
10
11      end for
12
13   end for
14  Compute class embedding $y=\tau(L)$;
15   Compute class distilled images ${\mathcal{S}}_{L}=\{{\mathcal{D}}\circ{\mathcal{U}}_{t}(z^{k}_{L},y)\mid k=1,...,K\}$;
16
17 end for
Result: distilled images ${\mathcal{S}}=\bigcup_{L\in{\mathcal{L}}}{\mathcal{S}}_{L}$
Algorithm 4 D4M [50]
Appendix F Proof of main Theorem 1
We first require a lemma that controls diffusions for compact measures. In particular, the score is (weakly) monotonic.
Lemma 1 ([4]).
Let $\mu\in{\mathcal{P}}(\mathbb{R}^{d})$ be a probability measure with compact support, say bounded by $R$. Let $g_{t}(x)=\frac{1}{(2\pi t)^{-d/2}}\exp(-\|x\|^{2}/2t)$ be the density of the standard normal distribution in $\mathbb{R}^{d}$. Then if $p_{t}$ is the density of $\mu*g_{t}$, it satisfies
$\langle x-y,\nabla\log p_{t}(x)-\nabla\log p_{t}(y)\rangle\leq\left(\frac{R^{2}}{t^{2}}-\frac{1}{t}\right)\|x-y\|^{2}.$
Proof.
From [4, Sec. 2.1], the density $p_{t}$ can be written as
$p_{t}(x)=\frac{1}{(2\pi t)^{d/2}}\exp\left(-\left(\frac{\|x\|^{2}}{2t}+W_{t}(x)\right)\right),$
where
$W_{t}(x)=-\log\int_{\mathbb{R}^{d}}\exp(\frac{\langle x,z\rangle}{t})\nu(\mathrm{d}z)-\log C_{\nu},$
with $C_{\nu}(x)=\int_{\mathbb{R}^{d}}\exp(-\|x\|^{2}/2t)\mu(\mathrm{d}x)$ and $\nu(\mathrm{dx})=C_{\nu}^{-1}\exp(-\|x\|^{2}/2t)\mu(\mathrm{dx})$. Moreover,
$0\leq-\nabla^{2}W_{t}\leq\frac{R^{2}}{t^{2}}I_{d}.$ (26)
Therefore,
$\log p_{t}(x)=-\frac{d}{2}\log(2\pi t)-\frac{\|x\|^{2}}{2t}-W_{t}(x)$
has Hessian satisfying
$\nabla^{2}\log p_{t}(x)\leq\left(\frac{R^{2}}{t^{2}}-\frac{1}{t}\right)I_{d}.$
Therefore, $\nabla\log p_{t}$ satisfies the desired monotonicity condition. Note that $R$ can be strengthened to $\frac{1}{2}\operatorname{diam}(\operatorname{supp}\mu)$. ∎
We next require the following proposition, which can be thought of as a stochastic version of Gronwall’s inequality. We present a simple version of the even stronger result available in [19].
Proposition 8 ([19, Lem. 3.6]).
Consider the diffusion
[…]
$\langle x-y,f(x,t)-f(y,t)\rangle+\frac{1}{2}\|g(x,t)-g(y,t)\|^{2}_{\mathrm{HS}}\leq\phi(t)\|x-y\|^{2}.$ (28)
Then for processes $X^{x}_{t},X^{y}_{t}$ starting at $x,y$ respectively under 27, it holds for all $t\in(0,T]$ that
[…]
Our goal is to “pushforward” the convergence of the Wasserstein distance from the latent space (of optimal quantizations) through the diffusion model (backwards SDE 9) into the image space/manifold. We proceed with the proof.
Theorem 5.
Consider the VESDE/Brownian motion
[…]
or the VPSDE/Ornstein–Uhlenbeck process
[…]
Then, for any initial data distribution $\mu\in{\mathcal{P}}_{2}(\mathbb{R}^{d})$ with compact support bounded by $R>0$, the assumptions for Proposition˜7 hold and the backwards diffusion process is well posed.
Suppose further that there are two distributions $\mu_{T},\nu_{T}$ at time $T$ that undergo the reverse diffusion process (with fixed initial reference measure $\mu$) up to time $t=\delta\in(0,T)$ to produce distributions $\mu_{\delta},\nu_{\delta}$. Then there exists a constant $C=C(\delta,R,d)\in(0,+\infty)$ such that if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ is an $L$-Lipschitz function, then the difference in expectation satisfies
$\|\mathbb{E}_{\mu_{\delta}}[f]-\mathbb{E}_{\nu_{\delta}}[f]\|\leq CL{\mathcal{W}}_{2}(\mu_{T},\nu_{T}).$ (32)
Proof.
The main idea is to push a Wasserstein-optimal coupling through the backwards SDE, then using Proposition˜8 and Proposition˜2. First fix a $\delta\in(0,T]$, and let $\operatorname{supp}\mu$ be bounded by $R$. Let $g_{t}=\frac{1}{(2\pi t)^{-d/2}}\exp(-\|x\|^{2}/2t)$ denote the distribution of the Gaussian $\mathcal{N}(0,tI_{d})$ in $d$ dimensions. By the Hörmander condition, the densities of a random variable $X_{t}$ with initial distribution $X_{0}\sim\mu$ undergoing the Brownian motion or Ornstein–Uhlenbeck process exist, and furthermore the forward and backward SDE processes are diffusions [29, 15, 39].
Step 1. (Monotonicity of the drifts.) For the Brownian motion,
$p_{t}(x)=(\mu(z)*g_{t})(x).$
For the Ornstein–Uhlenbeck process, the solution from initial condition $X_{0}=x_{0}$ is
$x_{t}=x_{0}e^{-t/2}+W_{1-e^{-t}}.$
The law of $X_{t}$ is thus $\mu(e^{t/2}x)*g_{1-\exp(-t)}$, where $\mu(e^{t/2}x)$ has support bounded by $Re^{-t/2}$.
Apply Lemma 1 with $\mu$ for the Brownian motion, and $\mu(e^{t/2}x)$ for the Ornstein–Uhlenbeck process. The corresponding backward SDEs (in forward time) for the Brownian motion and Ornstein–Uhlenbeck process as given by the forward time versions of 9 are
$\mathrm{d}x=\nabla\log p_{T-t}(x)\mathrm{d}t+\mathrm{d}W\quad\text{for Brownian motion;}$
$\mathrm{d}x=\left[\frac{x}{2}+\nabla\log p_{T-t}(x)\right]\mathrm{d}t+\mathrm{d}W\quad\text{for the OU process,}$
where $p_{T-t}(x)$ is the law of $X_{T-t}$ for $t\in[0,T)$.
Step 2. (Lipschitz w.r.t. initial condition after diffusion.) For the backwards SDEs, since the score is (weakly) monotone and the diffusion term is constant, the backwards SDEs satisfy the monotonicity condition in Proposition 8. Hence, there exists a constant $C$ such that for any $Y_{t}^{x},Y_{t}^{y}$ evolving according to the backwards SDE,
$\mathbb{E}\|Y_{T-\delta}^{x}-Y_{T-\delta}^{y}\|\leq C\|x-y\|.$ (33)
From the monotonicity condition and Proposition 8, the constants can be chosen to be as follows, noting the diffusion term is constant:
[…]
Step 3. (Lift to function expectation.) Now let $Y_{t},\hat{Y}_{t}$ be two diffusions, initialized with distributions $Y_{0}^{x}\sim\mu_{T}$,  $\hat{Y}_{t}^{x}\sim\nu_{T}$. Let $\gamma\in\Gamma(\mu_{T},\nu_{T})$ be any coupling. Define the “lifted coupling” $\hat{\gamma}$ on the measurable space $\left((\mathbb{R}^{d}\times\mathbb{R}^{d})\times\Omega,{\mathcal{B}}(\mathbb{R}^{d}\times\mathbb{R}^{d})\otimes{\mathcal{F}}\right)$, where $(\Omega,{\mathcal{F}},\mathbb{P})$ is the underlying (filtered) probability space of the diffusion, as the pushforward of the diffusion:
$\hat{\gamma}=(Y_{0}\mapsto Y_{T-\delta},\hat{Y}_{0}\mapsto\hat{Y}_{T-\delta},\iota_{\Omega})_{\sharp}(\gamma\otimes\mathbb{P})$ (35)
Marginalizing over $\mathbb{P}$, this is a (probability) measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ since the backward SDE paths are continuous: the backward SDEs admit the following integral formulation, where $h(r,Y_{r})$ is the drift term of the backward SDE:
[…]
Moreover, $\hat{\gamma}$ is a coupling between $Y_{T-\delta}\sim\mu_{\delta}$ and $\hat{Y}_{T-\delta}\sim\nu_{\delta}$. We compute:
[…]
The desired inequality follows by taking infimums over admissible couplings $\gamma\in\Gamma(\mu_{T},\nu_{T})$. ∎
Appendix G Approximate Timings
All times are done on Nvidia A6000 GPUs with 48GB of VRAM. We note that synthesis time as reported in [50, 51] do not include the time required to generate the latent variables, and thus are not sufficiently representative of the end-to-end time required to distill the dataset.
Table 5: Time required for each step of dataset distillation on ImageNet-1K. Synthesis requires application of the Stable Diffusion V1.5 model to each distilled latent variable, and soft label requires application of the pre-trained ResNet-18 model to each distilled image. Memory usage is constant between IPCs due to equal batch size.
Step Time (IPC 10) Time (IPC 100)
1 (Latent clustering) 8 hours 8 hours
2 (Synthesis) 2 hours 1 day
3 (Soft label) 1 hour 16 hours
4 (Training ResNet18) 2.5 hours 9 hours
Appendix H Experiment Hyperparameters
We detail the parameters when training the student networks from the distilled data. They are mostly similar to [50].
Table 6: Hyperparameter setting for ImageNet-1K experiments.
Setting Value
Network ResNet
Input size 224
Batch size 1024
Training epochs 300
Augmentation RandomResizedCrop
Min scale 0.08
Max scale 1
Temperature 20
Optimizer AdamW
Learning rate 2e-3 for Resnet18, 1e-3 otherwise
Weight decay 0.01
Learning rate schedule $\eta_{k+1}=\eta_{k}/4$ at epoch 250
Variance reduction (heuristic). We note that the variance of the number of cluster assignments can vary significantly, sometimes up to two orders of magnitude, such as in Figure˜1. After normalizing the cluster counts in Step 3 of Algorithm˜2 to give weights $w\in(0,1)$, most weights are very small, and do not contribute much to the neural network training. To reduce this effect, we use the per-centroid weights as follows,
$w_{k}^{(L)}=\sqrt{K}\sqrt{v_{k}^{(L)}\big{/}\sum_{j=1}^{K}v_{j}^{(L)}}.$ (36)
We note that other choices of variance reduction can also be used, and leave a more thorough search to future work.

### v3

@@
 Abstract
-Dataset distillation aims to find a synthetic training set such that training on the synthetic data achieves similar performance to training on real data, with orders of magnitude less computational requirements.
-Existing methods can be broadly categorized as either bi-level optimization problems that have neural network training heuristics as the lower level problem, or disentangled methods that bypass the bi-level optimization by matching distributions of data.
-The latter method has the major advantages of speed and scalability in terms of size of both training and distilled datasets.
-We demonstrate that when equipped with an encoder-decoder structure, the empirically successful disentangled methods can be reformulated as an optimal quantization problem, where a finite set of points is found to approximate the underlying probability measure by minimizing the expected projection distance.
-In particular, we link existing disentangled dataset distillation methods to the classical optimal quantization and Wasserstein barycenter problems, demonstrating consistency of distilled datasets for diffusion-based generative priors.
-We propose Dataset Distillation by Optimal Quantization, based on clustering in a latent space.
-Compared to the previous SOTA method D4M, we achieve better performance and inter-model generalization on the ImageNet-1K dataset with trivial additional computation, and SOTA performance in higher image-per-class settings.
-Using the distilled noise initializations in a stronger diffusion transformer model, we obtain SOTA distillation performance on ImageNet-1K and its subsets, outperforming diffusion guidance methods.
+Dataset distillation aims to find a small synthetic training set, such that training on the synthetic data achieves similar performance to training on a larger training dataset.
+Early methods solve this by interpreting the distillation problem as a bi-level optimization problem.
+On the other hand, disentangled methods bypass pixel-space optimization by matching data distributions and using generative techniques, leading to better computational complexity in terms of size of both training and distilled datasets.
+We demonstrate that by using latent spaces, the empirically successful disentangled methods can be reformulated as an optimal quantization problem, where a finite set of points is found to approximate the underlying probability measure.
+In particular, we link disentangled dataset distillation methods to the classical problem of optimal quantization, and are the first to demonstrate consistency of distilled datasets for diffusion-based generative priors.
+We propose Dataset Distillation by Optimal Quantization (DDOQ), based on clustering in the latent space of latent diffusion models.
+Compared to a similar clustering method D4M, we achieve better performance and inter-model generalization on the ImageNet-1K dataset using the same model and with trivial additional computation, achieving SOTA performance in higher image-per-class settings.
+Using the distilled noise initializations in a stronger diffusion transformer model, we obtain competitive or SOTA distillation performance on ImageNet-1K and its subsets, outperforming recent diffusion guidance methods.
 1 Introduction
 Training powerful neural networks requires a large amount of data, and thus induces high computational requirements.
-Dataset distillation (DD) targets this computational difficulty by changing the data, as opposed to other parts of training such as optimization or architecture [53].
+Dataset distillation (DD) targets this computational difficulty by changing the data, as opposed to other parts of training such as optimization or architecture (Wang et al., 2018).
 The DD objective consists of finding a synthetic training set, such that training a neural network on the synthetic data yields similar performance.
 There are several closely related notions of reducing computational load when training new models on datasets.
-Core-set methods find a subset of training data (as opposed to synthetic data) such that a model trained on said subset will have similar performance [30, 11].
-Model distillation, sometimes known as knowledge distillation, aims to train a smaller (student) model that predicts the output of a pre-trained (teacher) model [12, 40].
-Importance sampling methods accelerate training by weighting training data, finding examples that are more influential for training [36].
-For more detailed surveys on dataset distillation methods and techniques, we refer to [57, 44].
+Core-set methods find a subset of training data (as opposed to synthetic data) that achieve good training performance (Mirzasoleiman et al., 2020; Feldman, 2020).
+Model distillation, sometimes known as knowledge distillation, aims to train a smaller model that predicts the output of a larger model (Gou et al., 2021; Polino et al., 2018).
+Importance sampling methods accelerate training by weighting training data, finding examples that are more influential for training (Paul et al., 2021).
+For more detailed surveys on dataset distillation methods and techniques, we refer to (Yu et al., 2023; Sachdeva and McAuley, 2023).
+Training data distributionLow-dimensional latent spaceDistilled latent points and weightsDistilled data distribution $\{(\quad,\quad)\}$(1)Encoding(2)Clustering(3)Decoding
+Figure 1: Sketch of the proposed method pipeline.
+Using an encoder/decoder model, we map our high dimensional data to a low-dimensional space, which is then clustered using $k$-means.
+The clustered latent points and weights are then decoded to obtain the distilled data.
+This work argues that the weights are important when decoding; furthermore, “disentangled” distillation using an encoder-cluster-decoder framework is asymptotically consistent.
 1.1 Bi-level formulation of dataset distillation
-We first present the DD problem statement as given in [44].
-Denote a training set by $\mathcal{T}$ (more generally, distribution of training data), and the expected and empirical risks (test and training loss) by $\mathcal{R}$ and $\mathcal{L}$ respectively, evaluated for a network parameter $\Phi$.
-Given a learning algorithm $\Phi=\Phi_{S}$, which takes a dataset $S$ and returns an $\mathcal{L}$-optimal parameter, the goal of DD is to find a synthetic dataset $\mathcal{S}$ (of given size) minimizing the test loss discrepancy:
-$\mathcal{S}=\operatorname*{arg\,min}_{S}\left|\mathcal{R}(\Phi_{S})-\mathcal{R}(\Phi_{\mathcal{T}})\right|.$ (1)
+Denote a training set (more generally, distribution of training data) by $\mathcal{T}$, and the expected and empirical risks (test and training loss) by $\mathcal{R}$ and $\mathcal{L}$ respectively, evaluated for some parameter $\theta$.
+The goal of DD is to find a synthetic dataset $\mathcal{S}$ (of given size) minimizing the test loss discrepancy (Sachdeva and McAuley, 2023):
+$\mathcal{S}=\operatorname*{arg\,min}_{S}\left|\mathcal{R}(\operatorname*{arg\,min}_{\theta}{\mathcal{L}}(S))-\mathcal{R}(\operatorname*{arg\,min}_{\theta}{\mathcal{L}}({\mathcal{T}}))\right|.$ (1)
 This formulation is computationally intractable.
 Approximations include replacing the minimum discrepancy objective with maximum test performance, replacing the learning algorithm $\Phi$ with an inner neural network optimization problem, and solving the outer minimization problem using gradient methods.
-Common heuristic relaxations to the bi-level formulation 1 include meta-learning [53, 10], distribution matching [58], and trajectory matching [5].
-Other methods include neural feature matching [59, 28] and the corresponding neural tangent kernel methods [33, 32], representative matching [26], and group robustness [52].
-For better scaling, [5, 31] consider using generative priors such as GANs to generate more visually coherent images, replacing the need for optimization with a neural network inversion task and increasing performance.
+Common heuristic relaxations to the bi-level formulation 1 include meta-learning (Wang et al., 2018; Deng and Russakovsky, 2022), distribution matching (Zhao and Bilen, 2023), and trajectory matching (Cazenavette et al., 2022).
+Other methods include neural feature matching (Zhou et al., 2022; Loo et al., 2022) and the corresponding neural tangent kernel methods (Nguyen et al., 2021; 2020), representative matching (Liu et al., 2023b), and group robustness (Vahidian et al., 2024).
+For better scaling, Cazenavette et al.
+(2022); Moser et al.
+(2024) consider using generative priors such as GANs to generate more visually coherent images, increasing performance and replacing the need for optimization with a neural network inversion task.
 While the bi-level formulation follows naturally from the qualitative problem statement of dataset distillation, there are two main drawbacks, namely computational complexity and model architecture dependence.
-This limits the applicability to large-scale datasets and imposes additional requirements when using the distilled data.
-The first prototypical example of using a generative model to alleviate the aforementioned drawbacks of the bi-level formulation is by “factorizing” the high-dimensional image-space optimization.
-This consists of jointly training generative “hallucinator” networks to generate samples from a common set of learnable “base” points [25].
-This was able to increase cross-architecture performance when compared to previous image-space methods.
-1.2 Disentangled methods
-The dimensionality of the underlying bi-level optimization problems limits the applicability on large scale datasets, which are particularly useful for computationally limited applications.
-For example, the ImageNet-1K dataset consists of 1.2M training images, totalling over 120GB of memory [9].
+The dimensionality of the underlying optimization problems limit the applicability on large scale datasets, which are particularly useful for computationally limited applications.
+For example, the ImageNet-1K dataset consists of 1.2M training images, totalling over 120GB of memory (Deng et al., 2009).
 The full dataset ImageNet-21K consists of over 14M images and takes up around 1.2TB of memory, which is generally infeasible to train expert models on, and makes backpropagation through network training steps impossible.
-[56] is the first work to “disentangle” the bi-level optimization framework into three separate problems, named Squeeze, Recover and Relabel (SRe2L).
+1.2 Disentangled and diffusion methods
+Yin et al.
+(2023) is the first work to “disentangle” the bi-level optimization framework into three separate problems, named Squeeze, Recover and Relabel (SRe2L).
 In particular, the inner neural network optimization problem is replaced with matching statistics of batch-normalization layers.
-Curriculum Data Augmentation (CDA) uses adaptive training to get more performance [55].
-[24] considers optimizing images such that neural network features are close to Wasserstein barycenters of the training image features.
-[51] proposes Realistic Diverse and Efficient Dataset Distillation (RDED), which replaces the latent clustering objective with a patch-based adversarial objective.
-[50] considers clustering directly in the latent space of a latent diffusion model (LDM) [43], named Dataset Distillation via Disentangled Diffusion Model (D4M).
+Curriculum Data Augmentation (CDA) uses adaptive training to get more performance (Yin and Shen, 2024).
+Liu et al.
+(2023a) considers optimizing images such that neural network features are close to Wasserstein barycenters of the training image features.
+Sun et al.
+(2024) proposes Realistic Diverse and Efficient Dataset Distillation (RDED), which replaces the latent clustering objective with a patch-based adversarial objective.
+Su et al.
+(2024) considers clustering directly in the latent space of a latent diffusion model (LDM) (Rombach et al., 2022), named Dataset Distillation via Disentangled Diffusion Model (D4M).
 This avoids backpropagation when distilling and has constant memory usage with respect to images per class (IPC), a direct advantage over the linear memory scaling of optimization-based methods.
-Recent state-of-the-art methods consider fine-tuning diffusion models to have better statistics [14], and guiding the diffusion to maximize the influence of the distilled points [6], similarly to active learning.
-While dataset distillation has had extensive experimental effort, few proper theoretical justifications or computable formalizations exist in the literature.
-[44] proposes a high-level formulation based on minimizing the difference in test loss between learning on the full dataset versus the synthetic dataset.
-[22] considers dataset distillation as dependent on the desired inference task (typically classification with cross-entropy loss for image data).
-This allows them to interpret trajectory-matching as a mean-field control problem, and find analogous procedures for non-imaging examples like finding collocation points for training physics informed neural networks.
-However, they do not consider the problem of consistency or convergence of the distilled datasets, which we aim to address in this work.
-We interpret dataset distillation as the task of finding a surrogate dataset that minimizes the Wasserstein distance with the original training dataset.
-As a consequence for training neural networks, the gradient steps for a well-distilled dataset will be similar to the steps for the training dataset.
-This generalizes the objectives of the bi-level approaches to allow for different network architectures.
-1.3 Contributions
+Recent state-of-the-art methods consider fine-tuning diffusion models to have better statistics (Gu et al., 2024), and guiding the diffusion to maximize the influence of the distilled points (Chen et al., 2025), similarly to active learning.
+While dataset distillation has had extensive experimental effort, few theoretical justifications or computable formalizations exist in the literature.
+Sachdeva and McAuley (2023) proposes a high-level formulation based on minimizing the difference in test loss between learning on the full dataset versus the synthetic dataset.
+Kungurtsev et al.
+(2024) considers dataset distillation as dependent on the desired inference task (typically classification with cross-entropy loss for image data), interpreting trajectory matching as a mean-field control problem.
+No prior literature on dataset distillation addresses theoretically whether or not the distilled datasets are reasonable approximations of the input training data distribution.
+In this work, we address convergence in measure space for disentangled methods, demonstrating consistency and convergence of the distilled datasets.
 We summarize the contributions of this work as follows.
   1. 1.
-    Disentangled dataset distillation can be theoretically justified using classical notions of optimal quantization and Wasserstein distance.
-Motivated by the empirical usage of clustering in latent spaces, we show in Theorem˜1 that optimal quantizations induce convergent approximations of gradients of population risk, and furthermore the rate is given by $\mathcal{O}(K^{-1/d})$, where $d$ is the dimension of the latent space and $K$ is the number of quantization points.
-This motivates the usage of a low-dimensional latent space.
+    We theoretically justify the disentangled dataset distillation framework, exploiting its structure to show that these methods converge to the true data distribution as the number of distilled points increases, using classical notions of optimal quantization and Wasserstein distance.
+Motivated by the empirical usage of clustering in latent spaces, we show in Theorem˜1 that optimal quantizations induce convergent approximations of gradients of population risk.
+Furthermore, Corollary˜1 shows the approximation rate is given by $\mathcal{O}(K^{-1/d})$, where $d$ is the dimension of the latent space and $K$ is the number of quantization points.
+This motivates the usage of a low-dimensional latent space to model the data distribution.
   2. 2.
-    From the aforementioned classical literature, we propose Dataset Distillation by Optimal Quantization (DDOQ) in Algorithm˜2, based on clustering in a low-dimensional latent space.
-This is related to a recent disentangled method D4M.
-Our proposed method has a smaller Wasserstein distance between the distilled latent points and latent data distribution.
-This theoretically results in network gradients that more closely approximate those from training on the full training dataset.
+    We propose Dataset Distillation by Optimal Quantization (DDOQ) in Algorithm˜1, a DD algorithm based on clustering in a low-dimensional latent space.
+Compared to a recent disentangled method D4M, our proposed method has a smaller Wasserstein distance between the distilled latent points and latent data distribution, indicating better approximation.
   3. 3.
-    We algorithmically compare our proposed method with D4M and various baselines on the ImageNet-1K dataset using the same generative diffusion model backbone, demonstrating significantly better classification accuracy at varied IPC budgets and better cross-architecture generalization.
-Moreover, using the stronger diffusion transformer backbone, we provide a central comparison with SOTA disentangled and diffusion-based distillation methods, yielding competitive or better results than existing state-of-the-art on ImageNet-1K and some subsets.
-This paper is structured as follows.
+    We algorithmically compare our proposed method with D4M and various common disentangled baselines on the ImageNet-1K dataset using the same generative diffusion model backbone, demonstrating significantly better classification accuracy at varied IPC budgets and better cross-architecture generalization.
+To demonstrate the potential of DDOQ, we additionally use the stronger diffusion transformer (DiT) backbone, used in recent diffusion-based methods.
+We provide a central comparison with SOTA disentangled and diffusion-based distillation methods, yielding competitive or better results than existing SOTA methods on ImageNet-1K and its subsets.
+This work is structured as follows.
 Section˜2 covers related background and convergence results, including optimal quantization and diffusion models.
-Theorem˜1 in Section˜3 demonstrates consistency of the optimal quantizers when passed through diffusion-based generative priors to the image space, and proposes a simple extension of the D4M method by adding automatically-learned weights in the prototyping phase.
-We further fully explain the data distillation pipeline in a sequential and systematic manner from distillation, to training new models using the computed weights.
+Theorem˜1 in Section˜3 demonstrates consistency of the optimal quantizers when passed through diffusion-based generative priors to the image space, and motivates adding automatically-learned weights in the prototyping phase.
+Section˜3.1 details the data distillation pipeline in a sequential manner, from distillation to training new models using the computed weights.
 Section˜4 contains experiments of our proposed method against the previously SOTA D4M method as well as other recent SOTA baselines on the large-scale ImageNet-1K dataset.
 2 Background
-This section covers the necessary background and concepts required for the theoretical interpretation of DD.
-Section˜2.1 first covers necessary notation.
-Section˜2.2 introduces the optimal quantization problem, the concept of quadratic distortion, and links it to Wasserstein optimality.
-Section˜2.3 provides a short exposition of score-based diffusion models, covering well-posedness and classical divergence bounds.
-2.1 Definitions and Notation
 Define $\mathcal{P}_{2}(\mathbb{R}^{d})$ to be the set of probability measures on $\mathbb{R}^{d}$ with finite second moment, not necessarily admitting a density with respect to the Lebesgue measure.
-We use ${\mathcal{W}}_{2}$ to denote the Wasserstein-2 distance between two probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ [45], defined as
-${\mathcal{W}}_{2}(\mu,\nu)=\left(\inf_{\gamma\in\Gamma(\mu,\nu)}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\ \mathrm{d}\gamma(x,y)\right)^{1/2},$ (2)
-where $\Gamma(\mu,\nu)$ denotes the set of all couplings, i.e. joint probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$.
+We use ${\mathcal{W}}_{2}$ to denote the Wasserstein-2 distance between two probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ (Santambrogio, 2015).
 2.1 Optimal Quantization
-For a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with finite second moments, an optimal quantization (or vector quantization) at level $K$ is a set of points $\{x_{1},...,x_{K}\}\subset\mathbb{R}^{d}$ such that the $\mu$-averaged Euclidean distance to the quantized points is minimal.
+For a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, an optimal quantization (or vector quantization) at level $K$ is a set of points $\{x_{1},...,x_{K}\}\subset\mathbb{R}^{d}$ such that the $\mu$-averaged Euclidean distance to the quantized points is minimal.
 This can be formulated as the minimizer of the (quadratic) distortion, defined as follows.
@@
 An optimal quantization is a minimizer of $\mathcal{G}$.
-Note that the assumption that $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ has finite second moments implies that the quadratic distortion is finite for any set of points, and that an optimal quantization exists [34].
+Note that the assumption that $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ has finite second moments implies that the quadratic distortion is finite for any set of points, and that an optimal quantization exists (Pagès, 2015).
 We note that the optimal quantization weights are uniquely determined by the quantization points.
-This gives equivalence of the distortion minimization problem to the Wasserstein minimization problem, when restricted to measures of finite support [34].
+This gives equivalence of the distortion minimization problem to the Wasserstein minimization problem, when restricted to measures of finite support (Pagès, 2015).
 Proposition 1.
 Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$.
-Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial C_{i})=0$.
-Then the measure $\nu$ that minimizes the Wasserstein-2 distance 2 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$.
+Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi cells $\mu(\partial C_{i})=0$.
+Then the measure $\nu$ that minimizes the Wasserstein-2 distance 14 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$, where $\delta(x_{i})$ denotes the Dirac delta distribution at $x_{i}$.
 Moreover, the optimal coupling is given by the projection onto the centroids.
-For a given quantization $\mathbf{x}$} or quantization level $K$, the quadratic distortion satisfies respectively:
-${\mathcal{G}}({\mathbf{x}})^{1/2}=\inf\left\{{\mathcal{W}}_{2}(\nu,\mu)\mid\text{probability measures }\nu,\,\operatorname{supp}\nu\subset{\mathbf{x}}\right\};$ (5)
-$\operatorname*{arg\,min}_{{\mathbf{x}},\,|{\mathbf{x}}|=K}{\mathcal{G}}({\mathbf{x}})=\operatorname*{arg\,min}\left\{{\mathcal{W}}_{2}(\nu,\mu)\mid\text{probability measures }\nu,\,|\operatorname{supp}\nu|\leq K\ \right\}.$ (6)
-In other words, minimizing the quadratic distortion is equivalent to finding a Wasserstein-2 optimal approximation with a $K$-finitely supported (probability) measure.
+Proof.
+Deferred to Section˜C.2. ∎
+In other words, finding points that minimize the quadratic distortion is equivalent to finding a $K$-finitely supported (probability) measure, minimizing the Wasserstein-2 distance to the underlying measure.
 Remark 1.
-The case where the approximating measure is a uniform Dirac mixture is called the Wasserstein barycenter problem [8].
-The Wasserstein error of the Wasserstein barycenter is higher than that of the optimal quantization, but it admits an easily computable dual representation.
-Optimal quantizers converge in Wasserstein distance at a rate $\Theta(K^{-1/d})$ to their respective distributions as the number of particles $K$ increases [13].
-Moreover, any limit point of the optimal quantizers is an optimal quantizer of the limiting distribution [41].
-Moreover, it can be shown that if a sequence of distributions converges in Wasserstein distance $\mu_{n}\rightarrow\mu$, then so do the errors in quantization for a fixed quantization level [27, Thm.
-4].
-The quantizer can also be shown to have nice approximation properties when taking expectations of functions.
+The case where the approximating measure is a uniform Dirac mixture is called the Wasserstein barycenter problem (Cuturi and Doucet, 2014).
+The Wasserstein barycenter has higher error than the optimal quantization, but it admits an easily computable dual representation.
+The quantizer can be shown to have nice approximation properties when taking expectations of functions.
 A prototypical example for DD would have $f$ be the gradient of a neural network with respect to some loss function.
@@
 Proof.
-Deferred to Section˜C.1. ∎
+Deferred to Section˜C.3. ∎
 2.1.1 Solving the optimal quantization problem
-To find an optimal quantizer, we now introduce a commonly used quantization algorithm: the competitive learning vector quantization (CLVQ) algorithm [1].
-This is derived directly from gradient descent on the quadratic distortion, starting from an initialization in the convex hull of the support of $\mu$.
-For a given quantization, the gradient of the distortion has a representation in terms of $\mu$-centroids of the corresponding Voronoi cells, given explicitly in Section˜C.3.
-The CLVQ algorithm is presented in Algorithm˜1.
+To find an optimal quantizer, we use the competitive learning vector quantization (CLVQ) algorithm (Ahalt et al., 1990).
+This arises directly from gradient descent on the quadratic distortion.
+The gradient of the distortion has a representation in terms of $\mu$-centroids of the corresponding Voronoi cells, given explicitly in Section˜C.5.
+The CLVQ algorithm is presented in the appendix as Algorithm˜2.
 It consists of iterating: (i) sampling from $\mu$, (ii) computing the nearest cluster centroid, and (iii) updating the cluster centroid with weighted average.
-This algorithm is equivalent to the mini-batch $k$-means method when the stepsizes $\gamma_{i}$ are chosen to be the reciprocals of the number of points per-cluster [46, 37].
+CLVQ is equivalent to the mini-batch $k$-means method when the step-sizes $\gamma_{i}$ are chosen to be the reciprocals of the number of points per-cluster (Sculley, 2010; Pedregosa et al., 2011).
 We can thus interpret mini-batch $k$-means as finding a local minima of the optimal quantization problem.
-While this algorithm produces points ${\mathbf{x}}=(x_{1},...,x_{K})$, it remains to compute the associated weights approximating the measures of the Voronoi cells $\mu(C_{i})$ as in Proposition˜1.
-This can be done in an online manner within the same iterations [34].
-Data: initial cluster centers $x_{1}^{(0)},...,x_{K}^{(0)}$, step-sizes $(\gamma_{i})_{i\geq 0}$, $i\leftarrow 0$
-1 Initialize weights ${\mathbf{w}}=(w_{1},...,w_{K})=(1/K,...,1/K)$;
-2 while not converged do
-3   Sample $X_{i}\sim\mu$;
-4   Select “winner” $k_{\textrm{win}}\in\operatorname*{arg\,min}_{1\leq k\leq K}\|X_{i}-x_{k}^{(i)}\|$;
-5   Update $x_{k}^{(i+1)}\leftarrow(1-\gamma_{i})x_{k}^{(i)}+\gamma_{i}X_{i}$ if $k=k_{\mathrm{win}}$, otherwise $x_{k}^{(i+1)}\leftarrow x_{k}^{(i)}$ ;
-6   Update weights $w_{k}\leftarrow(1-\gamma_{i})w_{k}+\gamma_{i}\mathbf{1}_{k=k_{\mathrm{win}}}$;
-7 $i\leftarrow i+1$;
-8
-9 end while
-Result: quantization $\nu_{K}=\sum_{k=1}^{K}w_{i}\delta(x_{k}^{*})$
-Algorithm 1 CLVQ
-Proposition 3 ([3, Prop. 7]).
+The CLVQ algorithm produces points ${\mathbf{x}}=(x_{1},...,x_{K})$, but it remains to compute the associated weights approximating the measures of the Voronoi cells $\mu(C_{i})$ as in Proposition˜1.
+This can be done in an online manner within the same iterations (Pagès, 2015).
+Proposition 3 (Bally and Pagès 2003, Prop. 7).
 Assume that the measure $\mu\in{\mathcal{P}}_{2+\eta}(\mathbb{R}^{d})$ for some $\eta>0$, and that it assigns no mass to hyperplanes.
-Assume further that the grids ${\mathbf{x}}^{(t)}$ produced by CLVQ Algorithm˜1 converge to a stationary grid $\mathbf{x}$$^{*}$, i.e.
+Assume further that the grids ${\mathbf{x}}^{(t)}$ produced by CLVQ converge to a stationary grid $\mathbf{x}$$^{*}$, i.e.
 $\nabla{\mathcal{G}}({\mathbf{x}}^{*})=0$, and that the step-sizes satisfy $\sum\gamma_{k}=+\infty$ and $\sum\gamma_{k}^{1+\delta}<\infty$ for some $\delta>0$.
@@
 In the following section, we show that this reduced distortion carries over to the image space, which leads to better training fidelity as given using Proposition˜2.
-Remark 2.
-Convergence to an optimal quantizer is only guaranteed in the case of a log-concave distribution in one dimension [21, 27].
-However, in higher dimensions, convergence to a (not necessarily optimal) stationary grid is possible [35, 34].
 2.2 Score-based diffusion
 To connect the quantization error on the latent space with the quantization error in the image space, we need to consider properties of the latent-to-image process.
-In particular, we focus on score-based diffusion models, interpreted as discretizations of particular noising SDEs [49].
+In particular, we focus on score-based diffusion models, seen as discretizations of particular noising SDEs (Song et al., 2020).
 Consider the following SDE, where $W$ is a standard Wiener process:
 […]
-Then, where the density of $x$ at time $t>0$ is given by $p_{t}$, the reverse of the diffusion process is given by the reverse-time SDE [16],
+The reverse of the diffusion process is given by the reverse-time SDE (Haussmann and Pardoux, 1986), where the density of $x$ at time $t>0$ is given by $p_{t}$,
 […]
-where $\bar{W}$ is the standard Wiener process running in reverse time from time $T$ to 0.
-For an increasing noising schedule $\sigma(t)\in[0,+\infty)$ or noise-scale $\beta(t)\in[0,1)$, the variance-exploding SDE (VESDE, or Brownian motion) is given by
+and $\bar{W}$ is a reverse time Wiener process.
+For an increasing noising schedule $\sigma(t)$ or noise-scale $\beta(t)$, the variance-exploding SDE (VESDE, or Brownian motion) and variance-preserving SDE (VPSDE, or Ornstein–Uhlenbeck process) are given respectively by
 […]
-and the variance-preserving SDE (VPSDE, or Ornstein–Uhlenbeck process) is given by
-[…]
-The VPSDE and VESDE are commonly used when training diffusion models due to their simple discretization.
-Indeed, they both arise as continuous limits of earlier proposed diffusion models.
-Specifically, VPSDE corresponds to denoising score matching with Langevin dynamics [48], and VESDE to denoising diffusion probabilistic models [47, 17].
-Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with bounded support.
-[39] demonstrates that the backward SDE processes are indeed diffusions for the VPSDE and VESDE.
-Moreover, the diffusion satisfies a data-processing inequality, showing that the $f$-divergence after backwards diffusion is at most the $f$-divergence at marginal time $T$.
-We review this in Appendix˜D.
-Convergence in $f$-divergence necessarily requires absolute continuity of the compared marginal with respect to the underlying diffused distribution, which is equivalent to continuity with respect to the Lebesgue measure by the Hörmander condition.
-This rules out singular initializations such as empirical measures, as one might want to use for dataset distillation.
-To combat this, we use weak convergence with appropriately chosen test functions.
-The goal in question: given Wasserstein-2 convergence of the marginals $\nu_{T}^{(k)}\rightarrow\mu_{T}$, we wish to derive a bound on expectations $\mathbb{E}_{\nu_{\delta}^{(k)}}[f]\xrightarrow{?}\mathbb{E}_{\mu_{\delta}}[f]$, for some fixed $\delta\in(0,T)$ and $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying some regularity conditions.
+These SDEs are related to early diffusion models.
+Specifically, VPSDE corresponds to denoising score matching with Langevin dynamics (Song and Ermon, 2019), and VESDE to denoising diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020).
+Using this particular structure, we may obtain convergence results as seen in the next section.
+We note that by time-rescaling, we may assume without loss of generality that the noising schedule is linear $\sigma^{2}(t)=t$ or the noise-scale is constant $\beta(t)=1$.
+The goal in question: given Wasserstein-2 convergence of the marginals $\nu_{T}^{(k)}\rightarrow\mu_{T}$, we wish to derive a bound on expectations $\mathbb{E}_{\nu_{\delta}^{(k)}}[f]\xrightarrow{?}\mathbb{E}_{\mu_{\delta}}[f]$, for some small fixed $\delta\in(0,T)$ and $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying some regularity conditions.
+In other words, we wish to show that generative diffusion preserves closeness of data distributions.
 Such a bound would directly link to training neural networks with surrogate data, e.g.
@@
 Remark 2.
-Having $\delta=0$ may not be well defined because of non-smoothness and blowup of the score at time 0 for singular measures [39, 54].
+Having $\delta=0$ may not be well defined because of non-smoothness and blowup of the score at time 0 for singular measures (Pidstrigach, 2022; Yang et al., 2023).
+Remark 3.
+Working with weak convergence is necessary due to the singular empirical measures.
+Pidstrigach (2022) demonstrates that the backward SDE process satisfies a data-processing inequality, showing that the $f$-divergence after backwards diffusion is at most the $f$-divergence at marginal time $T$.
+However, $f$-divergences require absolute continuity of the compared marginal with respect to the underlying diffused distribution, which is equivalent to absolute continuity with respect to the Lebesgue measure by the Hörmander condition.
+This rules out singular initializations such as empirical measures, which arise in dataset distillation.
 3 Dataset Distillation as Optimal Quantization
 Our main result Theorem˜1 gives consistency of dataset distillation for a score-based diffusion prior in the image space.
-It transforms convergence of the terminal distribution into convergence of the backwards-diffused distributions at small time.
 Later, we use this in Section˜3.1 to present DDOQ as a modification of the D4M method.
 By simply changing the clustering objective from a Wasserstein barycenter to an optimal quantization by adding weights, we can effectively reduce the Wasserstein distance to the data distribution.
+Moreover, from Proposition˜3, the weights are automatically determined during the $k$-means clustering process.
 Theorem 1.
-Consider the VESDE/Brownian motion
-[…]
-or the VPSDE/Ornstein–Uhlenbeck process
+Consider the VESDE/Brownian motion or the VPSDE/Ornstein–Uhlenbeck process
 […]
@@
 This bypasses the heuristics needed in bi-level DD formulations, and avoids fine-tuning or generation-time guidance of the diffusion models.
-Theorem˜1 combined with the asymptotic rates of [13] gives convergence rates as the number of quantization points increases.
-We note that this can be further be combined with the convergence of optimal quantization rates of empirical measures such as Theorem˜3 in Section˜C.2.
+Theorem˜1 combined with the asymptotic rates of Graf and Luschgy (2000) gives convergence rates as the number of quantization points increases.
+We note that this can be further be combined with the convergence of optimal quantization rates of empirical measures such as Theorem˜3 in Section˜C.4.
+In particular, the next result shows that as the number of points increases, we have convergence to the underlying data distribution in image space, giving consistency.
 Corollary 1.
@@
 $\|\mathbb{E}_{\mu_{\delta}}[f]-\mathbb{E}_{\nu^{(K)}_{\delta}}[f]\|=L\mathcal{O}(K^{-1/d}).$ (10)
-In particular, as the number of points increases, we have convergence to the underlying data distribution in image space, giving consistency.
 Data: Training data and labels $({\mathcal{T}},{\mathcal{L}})$, pre-trained encoder-decoder pair $({\mathcal{E}},{\mathcal{D}})$, text encoder $\tau$, latent diffusion model ${\mathcal{U}}_{t}$, target IPC count $K$
@@
 Algorithm 1 Dataset Distillation by Optimal Quantization (DDOQ)
+Table 1: Wasserstein-2 distances of the distilled latents compared to the encoded training data on classes of ImageNet-1K. We observe the weighting drops the Wasserstein-2 distance significantly.
+Test class 0 1 2 Avg reduction
+IPC 10 D4M 49.67 51.14 47.16 -15.7%
+DDOQ 42.97 42.71 39.14
+IPC 50 D4M 47.34 49.10 45.28 -16.1%
+DDOQ 40.41 41.01 37.53
 3.1 Proposed method: Dataset Distillation by Optimal Quantization
 We have seen that clustering gives a consistent approximation to the latent distribution.
-Moreover, using the generative diffusion model, this then gives a consistent approximation in the image space.
-Now we present the proposed algorithm, Dataset Distillation by Optimal Quantization (DDOQ).
-We detail the steps in text below and summarize in Algorithm˜2.
-Suppose we are given an LDM, i.e.
-a (conditional) encoder-decoder pair and a diffusion-based model on the latent space [43].
+Using the generative diffusion model and Theorem˜1, we obtain a consistent approximation to the original data distribution in the image space.
+We now propose Dataset Distillation by Optimal Quantization (DDOQ).
+We detail the steps in text below and summarize in Algorithm˜1.
+Suppose we are given a latent diffusion model (LDM), i.e.
+a (conditional) encoder-decoder pair and a diffusion model on the latent space (Rombach et al., 2022).
 Let $K$ be the target number of images per class.
-Constructing the distilled dataset consists of the following steps in sequence.
+Constructing and using the distilled dataset consists of the following steps in sequence.
   1. 1.
@@
     $\min_{\theta}\sum_{(x,y,w)}w\cdot\ell(x,y,\theta).$
+Within Figure˜1, the encoder and decoder are given by a pre-trained encoder-decoder model, combined with a diffusion model in the latent space.
+Compared with D4M, we include (automatically determined) weights when training a new network in the final step, indicated by the variable length arrows.
+These are justified by considering the expectation of network gradients with respect to the training and distilled distributions.
+As seen in Table˜1, the inclusion of the weights significantly decreases the Wasserstein distance of the distilled points to the data distribution, when tested on classes of ImageNet-1K.
 We note that there is flexibility in the choice of label $y$ when training a new model.
@@
 4 Experiments
-We compare with two different latent diffusion architectures, namely the original latent diffusion model utilizing UNets [43], then the stronger diffusion transformer (DiT) architecture [38].
+We compare with two different latent diffusion architectures, namely the original latent diffusion model utilizing UNets (Rombach et al., 2022), then the stronger diffusion transformer (DiT) architecture (Peebles and Xie, 2023).
 We first use the LDM to compare the pure algorithmic differences of DDOQ with D4M and related baselines.
@@
 However, the weights are indicative of the distribution of the training data in the latent space of the diffusion model.
-Table 1: Comparison of top-1 classification performance on the ImageNet-1K dataset for baselines versus the proposed DDOQ method at various IPCs.
-We observe that DDOQ outperforms the previous SOTA method D4M, due to the addition of weights to the synthetic data.
+Table 2: Comparison of top-1 classification performance on the ImageNet-1K dataset at various IPCs.
+We observe that the proposed DDOQ outperforms the clustering-based method D4M, due to the addition of weights to the synthetic data.
 The maximum performance for all methods should be 69.8 as the soft labels are computed using a pre-trained ResNet-18 model.
-Figures for TESLA, SRe2L and D4M are taken directly from [50], RDED and CDA from [51, 55].
+In particular, we achieve a 30% reduction in error gap using ResNet-101 at IPC 200.
 […]
-To validate the proposed DDOQ algorithm, we directly compare with the previous state-of-the-art methods D4M [50] and RDED [51] on the ImageNet-1K dataset.
-The latent diffusion model chosen for latent generation and image synthesis is the publicly available pre-trained Stable Diffusion V1.5 model.
-We report results for SRe2L [56], RDED [51], CDA [55] as given.
-For low IPC, we also report the TESLA method, which is a SOTA bi-level method based on MTT [7].
-RDED achieves SOTA results for low IPC using its aggregated images and special training schedule, while CDA improves upon SRe2L using time-varying augmentation.
-For consistency and a more direct comparison with previous methods, we use the pre-trained PyTorch ResNet-18 model to compute the soft labels, using the same protocol as [50].
-After computing the soft labels using the pre-trained ResNet-18 model, we train new ResNet-18, ResNet-50 and ResNet-101 models to match the soft labels.
-The data augmentation is also identical, with the only difference being the addition of the weights to the training objective and minimal optimizer hyperparameter tuning for the new objective.
-We provide a direct comparison of the distilled data performance in Table˜1 for the IPCs $K\in\{10,50,100,200\}$.
-Our figures for DDOQ are averaged over five models trained on the same distilled data.
+To validate the proposed DDOQ algorithm, we directly compare with the previous state-of-the-art disentangled methods D4M (Su et al., 2024) and RDED (Sun et al., 2024) on the ImageNet-1K dataset.
+Baseline figures are reported as given in their respective works.
+For low IPC, we also report the TESLA method, which is a SOTA bi-level method based on MTT (Cui et al., 2023), but is unscalable past IPC=10.
+RDED achieves strong results for low IPC using its aggregated images and special training schedule, while CDA improves upon SRe2L using time-varying augmentation.
+For consistency, soft labels are computed using the pre-trained PyTorch ResNet-18 model, and new ResNet-{18,50,101} models are trained using the distilled data.
+We provide a direct comparison of the distilled data performance in Table˜2 for the IPCs $K\in\{10,50,100,200\}$.
+Variances for DDOQ are averaged over five models trained on the same distilled data.
 We observe that while RDED is very powerful in the low IPC setting due to the patch-based distilled images, which effectively gives the information of 4 (down-sampled) images in one training sample.
@@
 We observe that DDOQ has uniformly better cross-architecture generalization for convolutional teacher and student architectures, and slightly worse performance for student models using the transformer architecture Swin-T.
-Figures for D4M as reported in [50].
-Bold indicates best method.
 Teacher Network Student Network
@@
 DDOQ 28.5 24.1 29.3 36.0
-Table 3: Wasserstein-2 distances of the distilled latents compared to the encoded training data. We observe the weighting drops the Wasserstein-2 distance by around 20%.
-Test class 0 1 2
-IPC 10 D4M 49.67 51.14 47.16
-DDOQ 42.97 42.71 39.14
-IPC 50 D4M 47.34 49.10 45.28
-DDOQ 40.41 41.01 37.53
-Table˜2 shows the generalization performance of the distilled latent points.
-Different PyTorch pre-trained teacher networks are used to create the soft labels as in Step 3 of DDOQ, in particular MobileNet-V2 and Swin-T.
+Table˜3 shows the generalization performance of the distilled latent points.
+The PyTorch pre-trained MobileNet-V2 and Swin-T networks are used to create the soft labels.
 We then evaluate the distilled images and soft labels with three convolutional architectures and the Swin-T transformer architecture.
@@
 The slightly worse performance of DDOQ when using a transformer architecture for the student may be due to more precise hyperparameter tuning requirements.
-To illustrate the weights, Figure˜1 plots ten example images from the “jeep” class when distilled using $K=10$ IPC.
+To illustrate the weights, Figure˜2 plots ten example images from the “jeep” class when distilled using $K=10$ IPC.
 We observe that there is a very large variance in the weights $v_{k}^{(L)}\big/\sum_{j=1}^{K}v_{j}^{(L)}$}, indicating the presence of strong clustering in the latent space.
-Nonetheless, there is no qualitative evidence that the weights indicate “better or worse” training examples.
-Indeed, among other classes, there are synthetic images that do not lie in the appropriate class, but are still weighted highly.
-To consider the effect of weighting the centroids, we consider the class-wise Wasserstein distance between the centroids and encoded training data, with and without the weights.
-Table˜3 plots the Wasserstein distance between the distilled latent points and the encoded training dataset, for the first three classes of ImageNet-1K.
-The latent points are clustered with uniform weights (D4M), and with variable weights (DDOQ).
-We observe that adding weights reduces the Wasserstein distance uniformly by around 20%.
-This leads to better gradient faithfulness when training a neural network.
+Nonetheless, there is no qualitative evidence that the weights indicate “better or worse” training examples, rather indicating the structure in the latent space.
 4.2 DiT backbone
-To compare the potential of DDOQ, we use a stronger generative model, namely the diffusion transformer [38].
+To compare the potential of DDOQ, we use a stronger generative model, namely the diffusion transformer (DiT) (Peebles and Xie, 2023).
 This architecture achieves uniformly better sample quality compared to the LDM used in the previous section, in terms of Inception Score and Fréchet inception distance.
-We compare with the SOTA DD methods based on diffusion guidance, namely Minimax [14] and Influence Guided Diffusion (IGD) [6], which both use DiT.
+We denote the method with DiT backbone as DDOQ-DiT.
+We compare with the SOTA DD methods based on diffusion guidance, namely Minimax (Gu et al., 2024) and Influence Guided Diffusion (IGD) (Chen et al., 2025), which both use DiT.
 These methods guide the decoding process using some fine-tuning or batch statistics.
-In contrast, DDOQ directly modifies the initialization in the latent space.
-In addition, we compare with samples directly generated with DiT from random initializations.
+In contrast, DDOQ-DiT directly modifies the initialization in the latent space.
+In addition, we compare with samples directly generated with DiT from random initializations, labelled ‘DiT’.
 We compare on ImageNet-1K, as well as the 10-class subsets ImageNette and ImageWoof.
-The baseline figures are taken from [6] which report higher numbers than [14].
-Table˜4 demonstrates that we significantly outperform the diffusion-guidance based methods on the full ImageNet-1K, as well as DDOQ with the UNet backbone.
+The baseline figures are taken from Chen et al.
+(2025) which report higher numbers than Gu et al.
+(2024).
+Table˜4 demonstrates that we significantly outperform the diffusion-guidance based methods on the full ImageNet-1K, as well as DDOQ-DiT with the UNet backbone.
 Moreover, we are competitive-with or better than the baselines on the 10-class subsets, namely outperforming in the low IPC setting and in the more difficult ImageWoof subset.
+Moreover, the stronger diffusion model significantly benefits the low IPC ImageNet-1K setting, increasing ResNet-18 test accuracy from 33.1 to 53.0.
+An ablation on the effect of different weights during training is given in Appendix˜I.
 Table 4: Comparisons on ImageWoof and ImageNette using the ResNetAP-10 architecture, and ImageNet-1K using ResNet-18.
-We observe that DDOQ-DiT is able to outperform the guided diffusion methods at low IPC and are competitive at higher IPC, namely outperforming the compared methods on ImageNet-1K.
-The maximum accuracies are 87.2, 94.6 and 69.8 for ImageWoof, ImageNette and ImageNet-1K respectively.
+DDOQ-DiT outperforms the guided diffusion methods at low IPC and is competitive at higher IPC, namely outperforming the compared methods on ImageNet-1K.
 Dataset IPC DiT DiT-IGD Minimax Minimax-IGD DDOQ-DiT Full
@@
 Inspired by optimal quantization and Wasserstein theory, we theoretically demonstrate consistency of the distilled datasets in Theorem˜1 when using standard diffusion-based generative models to generate the synthetic data.
-We empirically demonstrate performance higher than previous state-of-the-art methods on a large-scale ImageNet-1K dataset, as well as in the low IPC setting of ImageNet subsets.
-We have presented theoretical justification for one framework of dataset distillation.
-Future work could include sharper bounds in Theorem˜1 that exploit the sub-Gaussianity of the diffused distributions or manifold hypothesis, or investigating alternative choices of weight variance reduction.
-Other interesting directions could be relating the weightings of the synthetic data to the hardness of learning the data, such as in [20], further increasing the performance using different training regimes such as curriculum learning, or extending the theoretical analysis to inexact score matching.
-Possible empirical work could investigate correlations between the weights and influence [42].
-Acknowledgments and Disclosure of Funding
+Experiments show the proposed method is competitive or better than SOTA on large-scale ImageNet experiments.
+We have presented theoretical justification for disentangled dataset distillation methods, which rely on clustering as the mechanism for approximating the underlying data distributions.
+More specifically, we justify the combination of a low-dimensional latent space and a diffusion model.
+Consistency or convergence of other dataset distillation frameworks such as bi-level methods and diffusion guidance are still open questions.
+Future work could include sharper bounds in Theorem˜1 that exploit the sub-Gaussianity of the diffused distributions or manifold hypothesis.
+Other interesting directions could be relating the weightings of the synthetic data to the hardness of learning the data, such as in (Joshi and Mirzasoleiman, 2023), further increasing the performance using different training regimes such as curriculum learning, or extending the theoretical analysis to inexact score matching.
+Possible empirical work could investigate correlations between the weights and influence (Pruthi et al., 2020).
+Acknowledgments
 This work was done while HYT was on placement in GSK.ai. HYT acknowledges support from GSK and the Masason Foundation.
@@
   1. 1.
-    (Meta-learning.) This uses the assumption that more data produces better results, replacing the risk matching objective with a risk minimization objective and removing the dependence on $\mathcal{T}$ [53, 10].
+    (Meta-learning.) This uses the assumption that more data produces better results, replacing the risk matching objective with a risk minimization objective and removing the dependence on $\mathcal{T}$ (Wang et al., 2018; Deng and Russakovsky, 2022).
 The learning algorithm $\Phi_{S}=\Phi_{S}(\theta_{0})$ with initial parameter $\theta_{0}$ is given by unrolling gradient steps $\theta_{t+1}=\theta_{t}-\eta\nabla\mathcal{L}_{S}(\theta_{t})$.
@@
   2. 2.
-    (Distribution matching.) Inspired by reproducing kernel Hilbert spaces and the empirical approximations using random neural network features, [58] proposes an alternate distillation objective that is independent of target loss.
+    (Distribution matching.) Inspired by reproducing kernel Hilbert spaces and the empirical approximations using random neural network features, Zhao and Bilen (2023) proposes an alternate distillation objective that is independent of target loss.
 The minimization objective is
@@
   3. 3.
-    (Trajectory matching.) For a fixed network architecture, this method aims to match the gradient information of the synthetic and true training datasets from different initializations [5].
+    (Trajectory matching.) For a fixed network architecture, this method aims to match the gradient information of the synthetic and true training datasets from different initializations (Cazenavette et al., 2022).
 The heuristic is that similar network parameters will give similar performance.
@@
     Practical computational approximations include pre-training a large number of teacher networks from different initializations.
-Appendix B Proof of Proposition 1
+Appendix B Additional definitions
+The Wasserstein-2 distance is defined as
+${\mathcal{W}}_{2}(\mu,\nu)=\left(\inf_{\gamma\in\Gamma(\mu,\nu)}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\ \mathrm{d}\gamma(x,y)\right)^{1/2},$ (14)
+where $\Gamma(\mu,\nu)$ denotes the set of all couplings, i.e. joint probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$.
+Appendix C Optimal quantization
+C.1 Additional background
+Optimal quantizers converge in Wasserstein distance at a rate $\Theta(K^{-1/d})$ to their respective distributions as the number of particles $K$ increases (Graf and Luschgy, 2000).
+Moreover, any limit point of the optimal quantizers is an optimal quantizer of the limiting distribution (Pollard, 1982).
+Moreover, it can be shown that if a sequence of distributions converges in Wasserstein distance $\mu_{n}\rightarrow\mu$, then so do the errors in quantization for a fixed quantization level (Liu and Pagès, 2020, Thm.
+4).
+C.2 Proof of Proposition 1
 Proposition 4.
@@
 Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial(C_{i}))=0$.
-Then the measure $\nu$ that minimizes the Wasserstein-2 distance 2 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$.
+Then the measure $\nu$ that minimizes the Wasserstein-2 distance 14 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$.
 Moreover, the optimal coupling is given by the projection onto the centroids.
@@
 from the null condition).
-This shows a lower bound of 2 that is attained.
+This shows a lower bound of 14 that is attained.
 ∎
-Appendix C Optimal quantization
-C.1 Proof of Thm 2
+C.3 Proof of Proposition 2
 Theorem 2.
@@
 The first equality comes from definition of $\nu$, and the inequalities from the Lipschitz condition and Hölder’s inequality respectively. ∎
-C.2 Convergence of optimal quantization
+For a given quantization $\mathbf{x}$} or quantization level $K$, the quadratic distortion thus satisfies respectively:
+${\mathcal{G}}({\mathbf{x}})^{1/2}=\inf\left\{{\mathcal{W}}_{2}(\nu,\mu)\mid\text{probability measures }\nu,\,\operatorname{supp}\nu\subset{\mathbf{x}}\right\};$ (16)
+$\operatorname*{arg\,min}_{{\mathbf{x}},\,|{\mathbf{x}}|=K}{\mathcal{G}}({\mathbf{x}})=\operatorname*{arg\,min}\left\{{\mathcal{W}}_{2}(\nu,\mu)\mid\text{probability measures }\nu,\,|\operatorname{supp}\nu|\leq K\ \right\}.$ (17)
+C.4 Convergence of optimal quantization
 The following result gives a convergence rate, assuming slightly higher regularity of the underlying probability measure.
-Proposition 5 ([27]).
+Proposition 5 (Liu and Pagès (2020)).
 Let $\eta>0$, and suppose $\mu\in{\mathcal{P}}_{2+\eta}(\mathbb{R}^{d})$. There exists a universal constant $C_{d,\eta}\in(0,+\infty)$ such that for every quantization level,
@@
 The following non-asymptotic result considers the convergence of optimal quantizers for a sequence of probability measures, converging in the Wasserstein sense.
-Theorem 3 ([27, Thm. 4]).
+Theorem 3 (Liu and Pagès 2020, Thm. 4).
 Fix a quantization level $K\geq 1$.
@@
 where $e^{*}_{K,\mu}=[\inf_{{\mathbf{x}}}{\mathcal{G}}_{K,\mu}({\mathbf{x}})]^{1/2}$ is the optimal error.
-C.3 Gradient interpretation of CLVQ
-Proposition 6 (Differentiability of distortion [34, Prop. 3.1]).
+Remark 4.
+Convergence to an optimal quantizer is only guaranteed in the case of a log-concave distribution in one dimension (Kieffer, 1982; Liu and Pagès, 2020).
+In higher dimensions, convergence to a stationary (but not optimal) grid is possible (Pages and Yu, 2016; Pagès, 2015).
+C.5 Gradient interpretation of CLVQ
+Proposition 6 (Differentiability of distortion Pagès 2015, Prop. 3.1).
 Let ${\mathbf{x}}=(x_{1},...,x_{K})\in(\mathbb{R}^{d})^{K}$ be such that the $x_{i}$ are pairwise distinct, and assume that $\mu(\partial C_{i})=0$.
@@
 where $\mathrm{Hull}$ denotes the convex hull.
-Recalling that the gradient 22 is a $\mu$-expectation over $C_{i}(x)$, the corresponding stochastic Monte Carlo version of the above gradient descent yields the CLVQ algorithm 1:
+Recalling that the gradient 20 is a $\mu$-expectation over $C_{i}(x)$, the corresponding stochastic Monte Carlo version of the above gradient descent yields the CLVQ algorithm 2:
 ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\gamma_{k}\left(\mathbf{1}_{X_{k}\in C_{i}^{(k)}}x_{i}^{(k)}-X_{k}\right)_{1\leq i\leq K},\quad X_{k}\sim\mu.$ (22)
 Note that the computation of this requires the ability to sample from $\mu$, as well as being able to compute the nearest neighbor of $X_{k}$ to the quantization set $\mathbf{x}$} (equivalent to the inclusion $X_{k}\in C_{i}^{(K)}$).
-C.4 Lloyd I algorithm
+Data: initial cluster centers $x_{1}^{(0)},...,x_{K}^{(0)}$, step-sizes $(\gamma_{i})_{i\geq 0}$, $i\leftarrow 0$
+1 Initialize weights ${\mathbf{w}}=(w_{1},...,w_{K})=(1/K,...,1/K)$;
+2 while not converged do
+3   Sample $X_{i}\sim\mu$;
+4   Select “winner” $k_{\textrm{win}}\in\operatorname*{arg\,min}_{1\leq k\leq K}\|X_{i}-x_{k}^{(i)}\|$;
+5   Update $x_{k}^{(i+1)}\leftarrow(1-\gamma_{i})x_{k}^{(i)}+\gamma_{i}X_{i}$ if $k=k_{\mathrm{win}}$, otherwise $x_{k}^{(i+1)}\leftarrow x_{k}^{(i)}$ ;
+6   Update weights $w_{k}\leftarrow(1-\gamma_{i})w_{k}+\gamma_{i}\mathbf{1}_{k=k_{\mathrm{win}}}$;
+7 $i\leftarrow i+1$;
+8
+9 end while
+Result: quantization $\nu_{K}=\sum_{k=1}^{K}w_{i}\delta(x_{k}^{*})$
+Algorithm 2 CLVQ
+C.6 Lloyd I algorithm
 Lloyd I.
@@
 Clearly, if this algorithm converges, then the centroids are equal to the $\mu$-centroids and the grid is stationary.
-This is more commonly known as the $k$-means clustering algorithm, employed in common numerical software packages such as scikit-learn [37].
+This is more commonly known as the $k$-means clustering algorithm, employed in common numerical software packages such as scikit-learn (Pedregosa et al., 2011).
 Convergence of the Lloyd-I algorithm can be found in e.g.
-[35].
+(Pages and Yu, 2016).
 Data: Probability distribution $\mu$ with finite first moment, initial cluster centers $x_{1}^{(0)},...,x_{K}^{(0)}$
@@
 Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}(\mathbb{R}^{d})$, assumed to have bounded support.
-Then, by the Hörmander condition, the law of a random variable $(X_{t})_{t\geq 0}$ evolving under either the VPSDE or the VESDE will admit a density $p_{t}(x)$ for all $t>0$ with respect to the Lebesgue measure, that is smooth with respect to both $x$ and $t$ [18].
-Using the following proposition, we have well-definedness of the backward SDE [2, 16].
+Then, by the Hörmander condition, the law of a random variable $(X_{t})_{t\geq 0}$ evolving under either the VPSDE or the VESDE will admit a density $p_{t}(x)$ for all $t>0$ with respect to the Lebesgue measure, that is smooth with respect to both $x$ and $t$ (Hörmander, 1967).
+Using the following proposition, we have well-definedness of the backward SDE (Anderson, 1982; Haussmann and Pardoux, 1986).
 Proposition 7.
-For the forward SDE 8, assume that there exists some $K>0$ such that
+For the forward SDE 5, assume that there exists some $K>0$ such that
   1. 1.
@@
   3. 3.
-    The solution $X_{t}$ of 8 has a $\mathcal{C}^{1}$ density $p_{t}(x)$ for all $t>0$, and
+    The solution $X_{t}$ of 5 has a $\mathcal{C}^{1}$ density $p_{t}(x)$ for all $t>0$, and
     $\int_{t_{0}}^{T}\int_{\|x\|<R}\|p_{t}\|^{2}+\|\nabla_{x}p_{t}(x)\|^{2}<+\infty\quad\forall t_{0}\in(0,T],\,R>0;$
@@
     The score $\nabla\log p_{t}(x)$ is locally Lipschitz on $(0,T]\times\mathbb{R}^{d}$.
-Then the reverse process $X_{T-t}$ is a solution of the 9, and moreover, the solutions of 9 are unique in law.
-Now given that the backwards SDE is indeed a diffusion, the data processing inequality uses the Markov property and states that the divergence after diffusion is less than the divergence before diffusion [23].
-This is summarized in [39, Thm.
-1].
+Then the reverse process $X_{T-t}$ is a solution of the 6, and moreover, the solutions of 6 are unique in law.
+Now given that the backwards SDE is indeed a diffusion, the data processing inequality uses the Markov property and states that the divergence after diffusion is less than the divergence before diffusion (Liese and Vajda, 2006).
+This is summarized in (Pidstrigach, 2022, Thm.
+1).
 Theorem 4.
-Denote the initial data distribution by $\mu_{0}=\mu$, and let $\mu_{T}$ be the distribution of a random variable $X_{t}$ satisfying the forward SDE 8 on $[0,T]$.
-Assume the above assumptions, and let $Y_{t}$ satisfy the backwards SDE 9 on $[0,T]$ with terminal condition $Y_{T}\sim\nu_{T}$.
+Denote the initial data distribution by $\mu_{0}=\mu$, and let $\mu_{T}$ be the distribution of a random variable $X_{t}$ satisfying the forward SDE 5 on $[0,T]$.
+Assume the above assumptions, and let $Y_{t}$ satisfy the backwards SDE 6 on $[0,T]$ with terminal condition $Y_{T}\sim\nu_{T}$.
 Denote by $\mu_{t}$ and $\nu_{t}$ the marginal distributions at time $t\in[0,T]$ of $X_{t}$ and $Y_{t}$, and assume that $\nu_{T}\ll\mu_{T}$.
@@
 Result: distilled images ${\mathcal{S}}=\bigcup_{L\in{\mathcal{L}}}{\mathcal{S}}_{L}$
-Algorithm 4 D4M [50]
+Algorithm 4 D4M (Su et al., 2024)
 Appendix F Proof of main Theorem 1
 We first require a lemma that controls diffusions for compact measures. In particular, the score is (weakly) monotonic.
-Lemma 1 ([4]).
+Lemma 1 (Bardet et al. (2018)).
 Let $\mu\in{\mathcal{P}}(\mathbb{R}^{d})$ be a probability measure with compact support, say bounded by $R$.
@@
 Proof.
-From [4, Sec. 2.1], the density $p_{t}$ can be written as
+From (Bardet et al., 2018, Sec. 2.1), the density $p_{t}$ can be written as
 $p_{t}(x)=\frac{1}{(2\pi t)^{d/2}}\exp\left(-\left(\frac{\|x\|^{2}}{2t}+W_{t}(x)\right)\right),$
@@
 Therefore, $\nabla\log p_{t}$ satisfies the desired monotonicity condition. Note that $R$ can be strengthened to $\frac{1}{2}\operatorname{diam}(\operatorname{supp}\mu)$. ∎
-We next require the following proposition, which can be thought of as a stochastic version of Gronwall’s inequality. We present a simple version of the even stronger result available in [19].
-Proposition 8 ([19, Lem. 3.6]).
+We next require the following proposition, which can be thought of as a stochastic version of Gronwall’s inequality.
+We present a simple version of the even stronger result available in Hudde et al.
+(2021).
+Proposition 8 (Hudde et al. 2021, Lem. 3.6).
 Consider the diffusion
@@
 $\langle x-y,f(x,t)-f(y,t)\rangle+\frac{1}{2}\|g(x,t)-g(y,t)\|^{2}_{\mathrm{HS}}\leq\phi(t)\|x-y\|^{2}.$ (26)
-Then for processes $X^{x}_{t},X^{y}_{t}$ starting at $x,y$ respectively under 27, it holds for all $t\in(0,T]$ that
+Then for processes $X^{x}_{t},X^{y}_{t}$ starting at $x,y$ respectively under 25, it holds for all $t\in(0,T]$ that
 […]
-Our goal is to “pushforward” the convergence of the Wasserstein distance from the latent space (of optimal quantizations) through the diffusion model (backwards SDE 9) into the image space/manifold.
+Our goal is to “pushforward” the convergence of the Wasserstein distance from the latent space (of optimal quantizations) through the diffusion model (backwards SDE 6) into the image space/manifold.
 We proceed with the proof.
@@
 Let $g_{t}=\frac{1}{(2\pi t)^{-d/2}}\exp(-\|x\|^{2}/2t)$ denote the distribution of the Gaussian $\mathcal{N}(0,tI_{d})$ in $d$ dimensions.
-By the Hörmander condition, the densities of a random variable $X_{t}$ with initial distribution $X_{0}\sim\mu$ undergoing the Brownian motion or Ornstein–Uhlenbeck process exist, and furthermore the forward and backward SDE processes are diffusions [29, 15, 39].
+By the Hörmander condition, the densities of a random variable $X_{t}$ with initial distribution $X_{0}\sim\mu$ undergoing the Brownian motion or Ornstein–Uhlenbeck process exist, and furthermore the forward and backward SDE processes are diffusions (Malliavin, 1978; Hairer, 2011; Pidstrigach, 2022).
 Step 1. (Monotonicity of the drifts.) For the Brownian motion,
@@
 Apply Lemma 1 with $\mu$ for the Brownian motion, and $\mu(e^{t/2}x)$ for the Ornstein–Uhlenbeck process.
-The corresponding backward SDEs (in forward time) for the Brownian motion and Ornstein–Uhlenbeck process as given by the forward time versions of 9 are
+The corresponding backward SDEs (in forward time) for the Brownian motion and Ornstein–Uhlenbeck process as given by the forward time versions of 6 are
 $\mathrm{d}x=\nabla\log p_{T-t}(x)\mathrm{d}t+\mathrm{d}W\quad\text{for Brownian motion;}$
@@
 All times are done on Nvidia A6000 GPUs with 48GB of VRAM.
-We note that synthesis time as reported in [50, 51] do not include the time required to generate the latent variables, and thus are not sufficiently representative of the end-to-end time required to distill the dataset.
+We note that synthesis time as reported in Su et al.
+(2024); Sun et al.
+(2024) do not include the time required to generate the latent variables, and thus are not sufficiently representative of the end-to-end time required to distill the dataset.
 Table 5: Time required for each step of dataset distillation on ImageNet-1K.
@@
 Appendix H Experiment Hyperparameters
-We detail the parameters when training the student networks from the distilled data. They are mostly similar to [50].
+We detail the parameters when training the student networks from the distilled data. They are mostly similar to Su et al. (2024).
+For consistency and a more direct comparison with previous methods, we use the pre-trained PyTorch ResNet-18 model to compute the soft labels, using the same protocol as Su et al.
+(2024).
+After computing the soft labels using the pre-trained ResNet-18 model, we train new ResNet-18, ResNet-50 and ResNet-101 models to match the soft labels.
+The data augmentation is also identical, so that the only differences are the addition of the weights to the training objective and some minor hyperparameter tuning for the new objective.
+The latent diffusion model chosen for latent generation and image synthesis is the publicly available pre-trained Stable Diffusion V1.5 model, the same as D4M.
 Table 6: Hyperparameter setting for ImageNet-1K experiments.
@@
 Variance reduction (heuristic).
-We note that the variance of the number of cluster assignments can vary significantly, sometimes up to two orders of magnitude, such as in Figure˜1.
-After normalizing the cluster counts in Step 3 of Algorithm˜2 to give weights $w\in(0,1)$, most weights are very small, and do not contribute much to the neural network training.
+We note that the variance of the number of cluster assignments can vary significantly, sometimes up to two orders of magnitude, such as in Figure˜2.
+After normalizing the cluster counts in Step 3 of Algorithm˜1 to give weights $w\in(0,1)$, most weights are very small, and do not contribute much to the neural network training.
 To reduce this effect, we use the per-centroid weights as follows,
 $w_{k}^{(L)}=\sqrt{K}\sqrt{v_{k}^{(L)}\big/\sum_{j=1}^{K}v_{j}^{(L)}}.$ (34)
-We note that other choices of variance reduction can also be used, and leave a more thorough search to future work.
+Appendix I Ablation on weighting
+To test the effect of the heuristic weights used in 34, we consider the setting of the DiT backbone Section˜4.2. We consider three different weightings within the training in Step 4:
+  1. 1.
+    Constant weights $w\equiv 1$. This is equivalent to D4M.
+  2. 2.
+    The heuristic weighting strategy 34.
+  3. 3.
+    The direct cluster weights, given by $w_{k}\propto v_{k}^{-1}$, where $v_{k}$ are the cluster counts from the $k$-means, normalized such that the sum over each class is 1.
+We use the same cluster images and only differ the weights.
+Table˜7 details the test accuracy for IPC 10 on ImageNet-1K with various weights, applied with a ResNet-18 teacher and ResNet-18 student architectures.
+We observe that the heuristic weighting strategy outperforms D4M across different learning rates.
+Moreover, the direct cluster weights are able to obtain higher test accuracy, albeit being more sensitive to the choice of learning rate.
+Table 7: Ablation across different weighting strategies when training a new student network with soft labels.
+We observe that the heuristic weight used in the main experiments always outperforms D4M.
+The direct cluster weighting is more sensitive to learning rate, but eventually outperforms the heuristic weighting for higher learning rates.
+Learning rate 1e-3 2e-3 5e-3 1e-2
+No weighting (D4M) 52.1 52.5 51.5 52.2
+Heuristic weight 34 52.3 52.6 53.2 53.6
+Direct cluster weighting 44.1 50.3 54.6 55.6